### v1

# Perfect Wave Transfer in Continuous Quantum Systems

Abstract
We study the perfect transfer of information in 1+1D continuous quantum systems. This includes effective descriptions of inhomogeneous spin chains, for which the notion of perfect state transfer in quantum information was introduced, and here phrased in terms of waves. We show that reflection symmetry is necessary for perfect wave transfer (PWT) in any inhomogeneous conformal field theory, and even sufficient when restricted to one-particle excitations. To determine if or when it is sufficient more generally, we first break conformal invariance and study a broad class of 1+1D bosonic theories. We show that the question can then be posed as an inverse Sturm-Liouville problem that determines when the bosonic theory exhibits PWT. We demonstrate how to uniquely solve this problem, which also shows that reflection symmetry is sufficient for the special case with conformal invariance. Using bosonization, our continuum results extend these notions to interacting quantum field theories.
Introduction.—Non-photonic quantum communication enables the efficient transmission of information from one quantum register to another [1]. This is desirable for quantum networks and computation [2, 3], not least since it allows for the perfect transfer of states even without dynamical control [4, 5]. First shown theoretically for communication channels in the form of a spin chain with spatially varying hopping, such perfect state transfer (PST) was extended to spins on graphs and other networks [6, 7, 8, 9, 10, 11, 12, 13, 14] and demonstrated experimentally [15, 16, 17, 18, 19, 20]. However, the notion has remained squarely in the discrete realm.
In this paper, we extend PST to continuous systems in one space and one time dimension (1+1D). In so doing, a rephrasing in terms of waves becomes natural, motivating the term perfect wave transfer (PWT). As for states in PST, a system on the interval $[-L/2,L/2]$ of length $L$ with suitable boundary conditions (BCs) exhibits PWT if the propagation of any initial wave, after some time $T$, amounts to its reflection; see Fig. 1. Concretely: The system enjoys PWT if
$\langle\mathcal{O}(x,0)\rangle=\langle\mathcal{O}(-x,T)\rangle,$ (1)
[…]
Related recent advances in the continuum include non-equilibrium studies of inhomogeneous versions of conformal field theory (CFT). These are conformally-invariant quantum field theories deformed so that the usual constant propagation velocity becomes a profile $v(x)$ [21, 22, 23, 24, 25, 26]. Among their applications, such inhomogeneous CFTs effectively describe gapless quantum many-body systems with mesoscopic inhomogeneities, such as spin chains with spatially varying hopping, trapped ultracold atoms, and arrays of quantum circuits; cf. [27]. Special cases include the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$ for some constant $v>0$ [28, 29]—the continuum counterpart of the inhomogeneous spin chain in [4] that first exhibited PST—and profiles produced by Möbius transformations [30, 31, 32, 33, 34, 35, 36, 37]. Note that for PWT, as for other transport properties of finite systems, the BCs [38, 39, 29, 40] are crucial. Furthermore, inhomogeneous CFT can be viewed as a theory in curved spacetime. As such, besides applications to quantum communication in solid-state devices [41, 42], our work also relates to relativistic quantum information [43, 44]. Since the same theories are also natural to study on the half or full infinite line, our results may likewise be relevant for long-distance quantum communication [45].
Figure 1: (a) Illustration of PWT with Neumann BCs showing an initial wave that evolves to its reflection at a certain time $T$. (b) Its manifestation in an inhomogeneous CFT with a given profile $v(x)/v_{0}$ (inset), showing a localized initial wave propagating along curved light-cone trajectories that recombine at the reflected point at $t=T\equiv L/v_{0}$ given by Eq. (7). Specifically, the magnitude of boson two-point correlations (17) at the free-fermion point is plotted, which by bosonization encodes the same information as $\log F(t)$ in Eq. (5) for delta-like waves $\xi_{1}$ centered at $-3L/8$ and $\xi_{2}$ at $x$ (setting $k_{\textrm{F}}=0$ for simplicity and $t^{\prime}=0$).
Engineering PST/PWT is at heart a spectral problem, amounting to whether the propagator coincides with the parity operator at certain times. We address this engineering question for quantum field theories describing a continuous quantum communication channel. Remarkably, for bosonic theories [46, 47, 48, 49, 50, 51, 52, 53], we arrive at an inverse problem for Sturm-Liouville operators, which we solve and whose solution determines if the channel enjoys PWT. Using bosonization [54, 55, 56, 57, 58, 59], our results extend the notion of PST/PWT to a large class of systems, including interacting fermionic quantum field theories.
Inhomogeneous CFT.—To begin with, we study the continuum limit of the inhomogeneous spin chains in [1, 4]. In the gapless regime, the continuous counterpart is an inhomogeneous CFT with a velocity profile $v(x)>0$, $x\in[-L/2,L/2]$ [21, 22, 23, 24, 25, 26]. The simplest example is that of inhomogeneous free fermions, given by the Hamiltonian (up to an overall constant and setting $\hbar=1$)
$H_{\textrm{iFF}}\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\frac{v(x)}{2}\Bigl{[}\!\left.:\!\psi^{\dagger}_{+}(-\mathrm{i})\partial_{x}\psi^{\vphantom{\dagger}}_{+}\!:\right.+\left.:\!\psi^{\dagger}_{-}\mathrm{i}\partial_{x}\psi^{\vphantom{\dagger}}_{-}\!:\right.+\mathrm{h.c.}\Bigr{]}$ (2)
[…]
We assume open (Neumann) BCs at $x=\pm L/2$. These can be incorporated by standard unfolding methods; cf. [61, 62, 57, 58, 38, 63]: Couple right and left movers by identifying $\psi_{-}(x)=\psi_{+}(L-x)$, giving a chiral theory of only right-moving fermions on $[-L/2,3L/2]$ with periodic BCs and velocity $v(x)$ for $x\in[-L/2,L/2]$ and $v(L-x)$ for $x\in[L/2,3L/2]$. This implies zero particle flow at $x=\pm L/2$ since the two fermion densities $\rho_{\pm}(x)=\left.:\!\psi^{\dagger}_{\pm}(x)\psi^{\vphantom{\dagger}}_{\pm}(x)\!:\right.$ are equal at each boundary and the current is $\propto\rho_{+}-\rho_{-}$.
PST was probed in [4] using a time-dependent overlap between one-particle states that are reflections of each other. Its continuum counterpart is to study states created by fields
$\psi(x)\equiv\mathrm{e}^{\mathrm{i}k_{\textrm{F}}x}\psi_{+}(x)+\mathrm{e}^{-\mathrm{i}k_{\textrm{F}}x}\psi_{-}(x),$ (3)
where $k_{\textrm{F}}$ denotes the Fermi momentum. Eq. (3) corresponds to the low-energy linearization of the spin-chain dispersion relation around its Fermi points ±$\pm k_{\textrm{F}}$, followed by recentering the resulting fields so they describe right/left-moving excitations around zero momentum. As a preparatory subset of possible states, consider
$|\Psi_{j}\rangle\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\xi_{j}(x)\psi^{\dagger}(x)|\Omega\rangle\qquad(j=1,2)$ (4)
for smooth enough functions $\xi_{j}(x)$, interpreted as distributions or “wave packets” of fermionic excitations in each state, and study the time-dependent overlap
$F(t)\equiv\langle\Psi_{2}|\mathrm{e}^{-\mathrm{i}H_{\textrm{iFF}}t}|\Psi_{1}\rangle.$ (5)
The states (4) form a restricted class as only one-particle excitations are possible and as right/left movers have the same particle distributions; cf. the more general treatment in the SM [60] where we lift the latter restriction.
To probe PWT for one-particle excitations as done in [4], we restrict to specular waves, $\xi_{2}(x)=\xi_{1}(-x)$, and compare $F(t)$ in Eq. (5) with $\langle\Psi_{1}|\Psi_{1}\rangle$. We say that: There is one-particle PWT if there exists $T>0$ so that
$|F(T)|=|\langle\Psi_{1}|\Psi_{1}\rangle|$ (6)
for all initial waves $\xi_{1}(x)$ in Eq. (4)  [^1] If Eq. (6) holds for any state $|\Psi_{1}\rangle$ in the Fock space, then general PST/PWT (i.e., for any number of excitations) would be enjoyed..
Eq. (6) amounts to a comparison between two-point correlation functions at equal and non-equal times. Such can be computed in any inhomogeneous CFT by mapping them to the corresponding homogeneous correlations [26]. The details are given in the SM [60] and boil down to comparing spacetime points after conformal transformations to coordinates given by
$y=y(x)\equiv\int_{0}^{x}\mathrm{d}s\,\frac{v_{0}}{v(s)},\qquad\frac{1}{v_{0}}\equiv\frac{1}{L}\int_{-L/2}^{L/2}\frac{\mathrm{d}s}{v(s)}.$ (7)
The outcome of the comparison shows that reflection symmetry is essential:
An inhomogeneous CFT exhibits one-particle PWT
if and only if $v(x)$ is even.
The corresponding times are odd multiples, $(2m+1)T$ for any integer $m$, of the time $T=L/v_{0}$ that excitations take to cross the entire system’s length. For later usage, we note that $y\in[-L/2,L/2]$ if $v(x)$ is even.
As an example, consider the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$, which gave the effective low-energy theory of the spin chain in [4]. This continuum theory exhibits PWT, just like for any even profile $v(x)$, in contrast to [4] where that profile was exceptional in this respect. Notably, the square-root profile is positive everywhere except at $x=\pm L/2$. Thus, strictly speaking, $v(x)$ should be regularized before using the tools in [26], but since the integral defining $1/v_{0}$ in Eq. (7) converges, it has no consequence for PWT. Oppositely, the limiting case of inhomogeneous CFT given by Möbius transformations so that $v(x)$ decays to zero at the boundaries cannot exhibit PWT since $1/v_{0}$ diverges.
Bosonic theories.—The above result and observations should not come as a surprise. After all, if one tries to compare discrete and continuous theories, then an inhomogeneous CFT can only describe low-energy excitations of a given lattice model. Moreover, one can show [60] that any Hamiltonian (with pure point spectrum) must commute with the parity operator for PWT to be possible, i.e., $v(x)$ must be even. The general structure of the spectrum due to conformal invariance in 1+1D [65] opens up the possibility that it might also be sufficient; e.g., the example (2) has a linear spectrum and is one-particle reducible, suggesting that if PWT holds for one-particle excitations then this should extend to many particles. The remainder of this paper concerns to what extent this is true more generally. Namely: What are necessary and sufficient conditions for PWT (1) for any number of excitations?
To answer this question, we first break conformal invariance. Examples include 1+1D fermions with non-local interactions [66, 67, 68], while here we focus on bosons, specifically inhomogeneous Tomonaga-Luttinger liquids (TLLs) [46, 47, 48, 49, 50, 51, 52, 53]. This theory is formulated using a compactified bosonic field $\varphi(x)$ (modulo $2\pi$) with conjugate $\Pi(x)$ satisfying $[\partial_{x}\varphi(x),\Pi(x^{\prime})]=\mathrm{i}\partial_{x}\delta(x-x^{\prime})$. The Hamiltonian is (up to an overall constant)
$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\biggl{(}\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}\biggr{)}\!:\right.\!,$ (8)
where, besides $v(x)>0$, the so-called Luttinger parameter $K(x)>0$ also depends on $x$. For simplicity, we assume that both functions are smooth. When $K(x)=K$ is constant, Eq. (8) is an inhomogeneous CFT, and when $K=1$, it is the bosonized version of the free fermions (2) [54, 55, 56, 57, 58, 59]. However, $K^{\prime}(x)\neq 0$ locally couples right and left movers non-trivially, thus they do not commute and the theory is not conformal [53]. Again, we impose open BCs, namely zero current at the boundaries:
$\jmath(x)\big{|}_{x=\pm L/2}=0,\quad\textnormal{where}\;\;\jmath(x)\equiv-\frac{v(x)K(x)}{\pi}\partial_{x}\varphi(x)$ (9)
is the total particle current. Unless stated otherwise, we also assume that $K^{\prime}(x)\big{|}_{x=\pm L/2}=0$ so that right and left movers are only coupled via the BCs (9) at those points.
Sturm-Liouville (SL) theory.—Following [69, 70, 71, 72, 73, 74, 27], we use SL theory to find eigenfunctions, using which the fields can be expanded and the Hamiltonian (8) diagonalized. (This generalizes [27] as $v(x),K(x)$ are unrelated and differs from the approach in [53].) Indeed, using integration by parts and Eq. (9),
$H_{\textrm{iTLL}}=\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}w(x)\!\left.:\!\Biggl{(}\biggl{[}\frac{\pi\Pi(x)}{w(x)}\biggr{]}^{2}+\varphi(x)\mathcal{A}\varphi(x)\Biggr{)}\!:\right.\!$ (10)
with the SL operator
$\mathcal{A}\equiv-\frac{1}{w(x)}\bigl{[}\partial_{x}p(x)\partial_{x}+q(x)\bigr{]},$ (11)
$w(x)\equiv\frac{K(x)}{v(x)},\quad p(x)\equiv v(x)K(x),\quad q(x)\equiv 0,$ (12)
where $\mathcal{A}$ acts on the space $\mathcal{D}$ of smooth  [^2] More precisely, twice continuously differentiable. functions $u(x)$ on $[-L/2,L/2]$ that satisfy $v(x)K(x)\partial_{x}u(x)\big{|}_{x=\pm L/2}=0$, i.e., Neumann BCs when $v(x),K(x)>0$.
We want to solve the SL problem
$\mathcal{A}u=\lambda u,\quad u\in\mathcal{D}.$ (13)
A distinction is warranted: Eq. (13) is regular if certain technical conditions are satisfied  [^3] Here, these amount to $v(x)$, $K(x)$ being strictly positive and $v(x)$, $K(x)$, $\partial_{x}p(x)$ [and $q(x)$ if non-zero] continuous on the whole interval $[-L/2,L/2]$.; otherwise it is irregular. Its importance stems from properties that hold for regular problems [77]:
  1. (i)
    The spectrum consists of eigenvalues $\lambda_{n}$ that are real and ordered: $0\leq\lambda_{0}<\lambda_{1}<\ldots$.
  2. (ii)
    Each $\lambda_{n}$ has a unique square-integrable eigenfunction $u_{n}(x)$ with $n$ isolated zeros in $[-L/2,L/2]$.
  3. (iii)
    The $u_{n}(x)$ form a complete orthonormal basis for square-integrable functions on $[-L/2,L/2]$ with weight $w(x)$: $\int_{-L/2}^{L/2}\mathrm{d}x\,w(x)u_{n}(x)u_{n^{\prime}}(x)=\delta_{n,n^{\prime}}$.
Furthermore, assuming that the integral defining $1/v_{0}$ in Eq. (7) converges:
  1. (iv)
    The $\lambda_{n}$ can at most grow as $n^{2}$ for large $n$, i.e., $\lambda_{n}=O(n^{2})$.
The last follows by Weyl’s law; see the SM [60] and, e.g., [78, 79]. Simply put, it gives a semi-classical estimate of the number of eigenvalues below a given large $\lambda>0$ for Laplace-type operators.
Unless otherwise stated, our SL problem will be regular. If irregular, the above properties are not guaranteed, while in certain cases they survive. Note also that, since $q(x)=0$ here, $u_{0}(x)=\textrm{const}$ solves Eq. (13) with $\lambda_{0}=0$.
The stated properties allow us to expand
$\varphi(x)=\sum_{n=0}^{\infty}\varphi_{n}u_{n}(x),\qquad\Pi(x)=\sum_{n=0}^{\infty}\Pi_{n}w(x)u_{n}(x),$ (14)
whereby we obtain $\bigl{[}\varphi_{n^{\vphantom{\prime}}},\Pi_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$ and can diagonalize $H_{\textrm{iTLL}}$ in Eq. (8). Indeed, the latter becomes
$H_{\textrm{iTLL}}=\frac{1}{2\pi}\bigl{[}\pi\Pi_{0}\bigr{]}^{2}+\sum_{n=1}^{\infty}E_{n}a^{\dagger}_{n}a^{\vphantom{\dagger}}_{n}$ (15)
with eigenenergies
$E_{n}=\sqrt{\lambda_{n}},$ (16)
where we expressed $\varphi_{n}=\sqrt{\pi/2E_{n}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right)$ and $\Pi_{n}=-\mathrm{i}\sqrt{E_{n}/2\pi}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ for $n\geq 1$ in bosonic operators $a^{\dagger}_{n}$, $a^{\vphantom{\dagger}}_{n}$ with $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$. Note that $\varphi_{0}$ (and zero modes in general) should be treated separately, but this will have no consequence here. In the SM [60], we also discuss SL operators (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq. (8).
The time evolution of any observable, such as $\varphi(x)$, will contain terms like $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$. Thus, as shown in the SM [60], to exhibit PWT as defined by Eq. (1), all $u_{n}(x)$ must have definite parity, i.e., be invariant under reflections $x\mapsto-x$ up to a sign, which must equal the value of each phase factor $\mathrm{e}^{\pm\mathrm{i}E_{n}T}$ for a common $T>0$. Moreover, the $u_{n}(x)$ have definite parity if and only if $v(x)$, $K(x)$ [and $q(x)$ if non-zero] are even, i.e., $H_{\textrm{iTLL}}$ is reflection symmetric. In conclusion, PWT boils down to requirements on the spectrum:
An inhomogeneous TLL exhibits PWT if and only if $v(x)$ and $K(x)$ are even and all eigenvalues are of the form $E_{n}=\pi m_{n}/T$ for certain integers $m_{n}$.
From Properties (i) and (ii) and since $\lambda_{0}=0$ here, these integers $m_{n}$ satisfy $0=m_{0}<m_{1}<\ldots$ with $m_{n}$ even (odd) if $n$ is even (odd). Moreover, from Property (iv), they scale as $m_{n}\sim n$ for large $n$, i.e., $m_{n}=n+o(n)$.
Lastly, to probe PWT based on Eq. (5), we note that the results above allow one to compute correlation functions for inhomogeneous TLLs; see the SM [60]. In particular, the ground-state $\varphi\varphi$ correlation function is
$\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle=\sum_{n=1}^{\infty}\frac{\pi}{2E_{n}}u_{n}(x)u_{n}(x^{\prime})\mathrm{e}^{-\mathrm{i}E_{n}(t-t^{\prime})}.$ (17)
Thus, for a spectrum satisfying the PWT requirements, $\langle\Omega|\varphi(x,T)\varphi(x^{\prime},0)|\Omega\rangle=\langle\Omega|\varphi(-x,0)\varphi(x^{\prime},0)|\Omega\rangle$, meaning that, for fixed $x^{\prime}$, the correlations at $t=T$ are reflections of those at $t=0$. In Fig. 1, we plot Eq. (17) for $K(x)=1$, which, by bosonization, is related to $\log F(t)$ in Eq. (5), in that it encodes the same information; see the SM [60].
Inverse Sturm-Liouville problem.—Reversing the logic, the PWT requirements on the spectrum yield an inverse spectral problem for SL operators with Neumann BCs [and $q(x)=0$]:
Given a spectrum $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ as above, determine the even functions $v(x)$ and $K(x)$.
A related prominent problem is the inverse scattering occurring in the theory of integrable PDEs, whose solution is given by the Gelfand-Levitan-Marchenko equation. However, that case is formulated on an infinite line, while here we are interested in a finite interval.
Before a general discussion, we consider two examples:
The inverse SL problem has a known solution when the spectrum is exactly linear: For $K(x)=K>0$ constant but $v(x)>0$ general, the SL operator is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq. (7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl{(}n\pi[L+2y]/2L\bigr{)}$ and $u_{0}(x)=\sqrt{v_{0}/L}$. This is precisely the conformal case of non-interacting right/left movers, enjoying PWT for any even $v(x)>0$.
In the previous example, the SL operator was regular. Relaxing this condition by taking $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl{[}1-(2x/L)^{2}\bigr{]}^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$, we get an irregular problem for which Properties (i)(iv) remain valid; the BCs being that a solution is sufficiently well-behaved for Eq. (9) to hold. The eigenfunctions are given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$ [80, §18.3] up to normalizations, which have definite parity and satisfy Eq. (13) with $\lambda_{n}=(\pi v_{0}/L)^{2}n(n+2\alpha)$ for $v_{0}=2v/\pi$. Clearly, only the case $\alpha=0$ (Chebyshev polynomials of the first kind) exhibits PWT, while all $\alpha\neq 0$ do not, including Legendre polynomials ($\alpha=1/2$) and Chebyshev polynomials of the second kind ($\alpha=1$) [80, §18.7]. The respective $\varphi\varphi$ correlations are plotted in Fig. 2.
To solve the inverse SL problem in general, we write the operator (11) in Liouville normal form, as is always possible for regular problems [80, §1.13(viii)]: Let $\mathfrak{u}(y)\equiv u(x)\sqrt{K(x)}$ for $y=y(x)$ given by Eq. (7) and define
Figure 2: Plot of $C(x,t)\equiv\operatorname{Re}\langle\Omega|\varphi(x,t)\varphi(-3L/8,0)|\Omega\rangle$ in Eq. (17) for eigenfunctions in the form of (a) Chebyshev polynomials of the first kind, which exhibit PWT, contrasted to (b) Legendre polynomials and (c) Chebyshev polynomials of the second kind, which do not. The upper row contains intensity plots of the correlations as functions of $x/L\in[-1/2,1/2]$ and $t/T\in[0,1]$, while the lower row compares their profiles at $t=0$ and $T$. Note the asymmetry in (b) and (c). The velocity $v(x)$ (same in all cases) and the respective Luttinger parameter $K(x)$ are shown in (d).
$\hat{q}(y)\equiv-\frac{q(x)}{w(x)}+\frac{1}{\sqrt{K(x)}}v_{0}^{2}\partial_{y}^{2}\sqrt{K(x)}.$ (18)
The eigenvalue problem (13) then becomes
$v_{0}^{2}\partial_{y}^{2}\mathfrak{u}(y)+[\lambda-\hat{q}(y)]\mathfrak{u}(y)=0$ (19)
with unchanged BCs  [^4] For even $v(x)>0$, they are $v_{0}\partial_{y}\mathfrak{u}(y)-v(x)\Lambda(x)\mathfrak{u}(y)=0$ at $y=\pm L/2$ with $\Lambda(x)\equiv\partial_{x}\ln\sqrt{K(x)}$ [53], which imply Neumann BCs if $K^{\prime}(x)\big{|}_{x=\pm L/2}=0$. and if reflection symmetric in $x$, then so in $y$. Clearly, $\mathfrak{u}_{n}(y)$ have the same eigenvalues as the respective $u_{n}(x)$ and similarly form a complete orthonormal basis but with identity weight.
Results in the literature often require two sets of eigenvalues for different BCs to get a unique solution to the inverse problem. However, one set suffices for reflection-symmetric SL operators and certain BCs, including ours. Indeed, [82] and [83, Theorem II] say that an even $\hat{q}(y)$ is uniquely determined by the spectrum for Neumann BCs. Thus, viewing $\hat{q}(y)$ as given and since $q(x)=0$, Eq. (18) becomes a second-order differential equation for $\sqrt{K(x(y))}$. By general uniqueness results [assuming $\hat{q}(y)$ is continuous] and since $K(x)$ must be even, this determines $K(x)$ up to changing coordinates (7) and an overall rescaling (discarding solutions that are not positive).
The results [82, 83] do not reconstruct $\hat{q}(y)$. However, since we know that $K(x)=K$ gives an SL problem with $m_{n}=n$ and $\hat{q}(y)=0$, a constant Luttinger parameter is the unique solution for the inverse problem given an exactly linear spectrum. Beyond this special case, nothing a priori prevents the spectrum from deviating from the leading $n$ behavior as long as the requirements on the $m_{n}$ remain satisfied. This would then give rise to a class of possible bosonic theories that exhibit PWT, where, by uniqueness, $\hat{q}(y)$ is non-zero, thus $K(x)$ non-constant. While not usually analytically tractable, there are constructive approaches to find $\hat{q}(y)$ [84, 85, 86].
Discussion.—We introduced the notion of PWT and showed that it is exhibited by one-particle excitations in inhomogeneous CFT whenever the velocity profile $v(x)>0$ is even. Breaking conformal invariance, by considering inhomogeneous TLLs with a Luttinger-parameter profile $K(x)>0$, we showed that PWT for general excitations amounts to an inverse spectral problem, which we solved. This continuum result for bosons is remarkably similar to earlier results: For spin chains with nearest neighbor hopping, the corresponding class of operators can be represented as Jacobi matrices, and the resulting inverse eigenvalue problem was studied [8, 87]. As in the continuum, two sets of eigenvalues are generally needed for uniqueness [88, 86], while one set suffices for matrices symmetric with respect to the anti-diagonal (corresponding to reflection symmetry) [89, Theorem 3].
By bosonization, our results for inhomogeneous TLLs cover fermionic theories with position-dependent interactions [53]. Indeed, even when $K(x)=K$ is constant, as long as $K\neq 1$, the theory describes an interacting quantum many-body system, albeit at low-energies.
Lastly, we showed that an inhomogeneous TLL with constant $K(x)=K$ enjoys PWT for general excitations whenever $v(x)>0$ is even. Such a theory is an example of an inhomogeneous CFT. It would be interesting to study PWT for any number of excitations in general CFTs, and also extend to systems on the half or full infinite line.
Acknowledgments.
Acknowledgements.
We are thankful to Pawel Caputa and Shinsei Ryu for motivating discussions on related topics. P.M. gratefully acknowledges financial support from the Wenner-Gren Foundations (Grant No. FT2022-0002). M.C. acknowledges financial support from the European Research Council (ERC Grant Agreement No. 818761), VILLUM FONDEN via the QMATH Centre of Excellence (Grant No. 10059) and the Novo Nordisk Foundation (Grant No. NNF20OC0059939 ‘Quantum for Life’), and thanks the NCCR SwissMAP of the Swiss National Science Foundation and the Section of Mathematics at the University of Geneva for their hospitality.
Supplemental Material
This supplemental material has three parts. Part A gives details for two-point correlations in inhomogeneous 1+1D CFT with open (Neumann) boundary conditions (BCs). Part B describes how to diagonalize general inhomogeneous Tomonaga-Luttinger liquids (TLLs) using Sturm-Liouville (SL) theory, leading to a formulation of perfect wave transfer (PWT) as an inverse SL problem, including extension to massive TLLs. Part C contains details for correlation functions in inhomogeneous TLLs.
Part A: Correlations in inhomogeneous 1+1D CFT with open boundaries
The general inhomogeneous CFT Hamiltonian is (setting $\hbar=1$)
$H_{\textrm{iCFT}}\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,v(x)\bigl{[}T_{+}(x)+T_{-}(x)\bigr{]},$ (20)
where $T_{\pm}(x)$ are the right/left-moving components of the stress-energy tensor in light-cone coordinates, whose Fourier transforms yield two commuting copies of the Virasoro algebra. In position space, they satisfy the equal-time commutation relations
$\bigl{[}T_{\pm}(x),T_{\pm}(x^{\prime})\bigr{]}=\mp 2\mathrm{i}\partial_{x}\delta(x-x^{\prime})T_{\pm}(x^{\prime})\pm\mathrm{i}\delta(x-x^{\prime})\partial_{x^{\prime}}T_{\pm}(x^{\prime})\pm\frac{c}{24\pi}\mathrm{i}\partial_{x}^{3}\delta(x-x^{\prime})$ (21)
and $\bigl{[}T_{\pm}(x),T_{\mp}(x^{\prime})\bigr{]}=0$. Note that, at non-equal time, $T_{\pm}$ should be understood to only depend on the “right/left-moving coordinate” $x^{\mp}$; for now, we only view these as labels, distinguishing dependencies on right/left components.
Besides the $T_{\pm}(x)$ fields, an important class of operators are Virasoro primary fields $\Phi(x^{-},x^{+})$, defined by how they behave under conformal transformations. The latter are given by orientation-preserving circle diffeomorphisms $f_{+}$, $f_{-}$ and primary fields transform as
$\Phi(x^{-},x^{+})\to f^{\prime}_{+}(x^{-})^{\Delta_{\Phi}^{+}}f^{\prime}_{-}(x^{+})^{\Delta_{\Phi}^{-}}\Phi(f_{+}(x^{-}),f_{-}(x^{+})),$ (22)
where ($\Delta_{\Phi}^{+},\Delta_{\Phi}^{-}$) denote the field’s conformal weights. To impose open BCs at $x=\pm L/2$, as in the main text, we will assume that primary fields come in pairs $\Phi_{+}$, $\Phi_{-}$ with conformal weights satisfying $\Delta^{+}_{\Phi_{+}}=\Delta^{-}_{\Phi_{-}}$, $\Delta^{-}_{\Phi_{+}}=\Delta^{+}_{\Phi_{-}}$, since this allows one to couple excitations so that right movers are reflected to left movers when reaching $x=L/2$, and vice-versa at $x=-L/2$. To ease notation, we will use $\Delta^{+}_{\Phi}\equiv\Delta^{+}_{\Phi_{+}}$, $\Delta^{-}_{\Phi}\equiv\Delta^{-}_{\Phi_{+}}$ for each such pair.
Before continuing, we mention two examples:
  1. Example 1:
    For the inhomogeneous free fermions in the main text, the components of the stress-energy tensor are $T_{\pm}(x)=(1/2)\bigl{[}\!\left.:\!\psi^{\dagger}_{\pm}(x)(\mp\mathrm{i}\partial_{x})\psi^{\vphantom{\dagger}}_{\pm}(x)\!:\right.+\mathrm{h.c.}\bigr{]}-\pi\big{/}12L^{2}$, and the primary fields of the theory are the fermionic fields $\psi_{\pm}(x)$ themselves, whose conformal weights are $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=1/2$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=0$.
  2. Example 2:
    For an inhomogeneous TLL with constant Luttinger parameter $K(x)=K$, one has $T_{\pm}(x)=({1}/{4\pi K})\!\left.:\!\bigl{[}\pi\Pi(x)\mp K\partial_{x}\varphi(x)\bigr{]}^{2}\!:\right.-\pi\big{/}12L^{2}$, and the primary fields include vertex operators constructed as exponentials of the bosonic fields. Applied to the Luttinger model [46, 47, 48], these vertex operators are (renormalized) fermionic fields $\psi_{\pm}$ with conformal weights $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=(1+K)^{2}/8K$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=(1-K)^{2}/8K$. (Clearly, free fermions are recovered for $K=1$.)
As in the main text, open BCs can be incorporated by standard unfolding methods; cf. [61, 62, 57, 58, 38, 63]: Couple right- and left-moving excitations by identifying
$T_{-}(x)=T_{+}(L-x),\qquad\Phi_{-}(x^{-},x^{+})=\Phi_{+}(L-x^{+},L-x^{-})$ (23)
for all pairs $\Phi_{+},\Phi_{-}$ of Virasoro primary fields. It follows that we can equivalently consider a chiral inhomogeneous CFT of only right movers on the interval $[-L/2,3L/2]$ of length $2L$ with velocity
$\underline{v}(x)\equiv\begin{cases}v(x)&\text{for }x\in[-L/2,L/2],\\ v(L-x)&\text{for }x\in[L/2,3L/2]\end{cases}$ (24)
and with every quantity satisfying periodic BCs. Note that the Hamiltonian (20) can then be written as
$H_{\textrm{iCFT}}\equiv\int_{-L/2}^{3L/2}\mathrm{d}x\,\underline{v}(x)T_{+}(x).$ (25)
For later reference, we define
$\underline{f}(x)\equiv\int_{0}^{x}\mathrm{d}s\,\frac{\underline{v}_{0}}{\underline{v}(s)},\qquad\frac{1}{\underline{v}_{0}}\equiv\frac{1}{2L}\int_{-L/2}^{3L/2}\frac{\mathrm{d}s}{\underline{v}(s)},\qquad\underline{x}_{t}^{\pm}(x)\equiv\underline{f}^{-1}\bigl{(}\underline{f}(x)\pm\underline{v}_{0}t\bigr{)}.$ (26)
Naturally, $\underline{f}$ lies in the universal cover $\widetilde{\operatorname{Diff}}_{+}(S^{1}_{2L})$ of the group of orientation-preserving diffeomorphisms of the circle $S^{1}_{2L}$ with circumference $2L$. Given any $v(x)$, one can show the following properties:
$\underline{v}(L-x)=\underline{v}(x),\qquad\underline{f}(L-x)=\underline{f}(L)-\underline{f}(x),\qquad\underline{f}^{-1}\bigl{(}\underline{f}(L)-y\bigr{)}=L-\underline{f}^{-1}(y),\qquad\underline{x}_{t}^{\pm}(L-x)=L-\underline{x}_{t}^{\mp}(x).$ (27)
Lastly, we introduce the undeformed Hamiltonian
$H_{0}=\int_{-L/2}^{3L/2}\mathrm{d}x\,\underline{v}_{0}T_{+}(x),$ (28)
whose ground state we denote by $|\Omega_{0}\rangle$. Similarly, we let $|\Omega\rangle$ denote the ground state of $H_{\textrm{iCFT}}$. Following [24, 26], the latter are related to their undeformed counterparts by
$U_{+}(\underline{f})H_{\textrm{iCFT}}U_{+}(\underline{f})^{-1}=H_{0}+\textrm{const},\qquad|\Omega\rangle=U_{+}(\underline{f})^{-1}|\Omega_{0}\rangle,$ (29)
where $U_{+}(\underline{f})$ is a projective unitary representation of $\underline{f}$ on the Hilbert space of the CFT.
Similar to the main text, we define
$|\Phi_{j}\rangle\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\left(\mathrm{e}^{-\mathrm{i}kx}\xi^{+}_{j}(x)\Phi_{+}^{\dagger}(x,x)+\mathrm{e}^{\mathrm{i}kx}\xi^{-}_{j}(x)\Phi_{-}^{\dagger}(x,x)\right)|\Omega\rangle\qquad(j=1,2)$ (30)
for smooth enough functions $\xi^{\pm}_{j}(x)$, interpreted as distributions or “wave packets” of right/left-moving excitations in each state, and some $k\in\mathbb{R}$ in order to accommodate for $k_{\textrm{F}}$ in Eq. (3). (The latter will have no effect on PWT, and one might as well set $k=0$.) As in the main text, this is a restricted class of states which only allow for one-particle excitations, while we eased other restrictions in the sense that the excitations are not necessarily fermionic and the distributions of right/left movers can be different.
To probe PWT for one-particle excitations, we seek to compute
$\langle\Phi_{j}|\Phi_{j}\rangle=\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\langle\Omega|\left(\mathrm{e}^{\mathrm{i}kx_{2}}\overline{\xi^{+}_{j}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(x_{2},x_{2})+\mathrm{e}^{-\mathrm{i}kx_{2}}\overline{\xi^{-}_{j}(x_{2})}\Phi_{-}^{\vphantom{\dagger}}(x_{2},x_{2})\right)\\ \times\left(\mathrm{e}^{-\mathrm{i}kx_{1}}\xi^{+}_{j}(x_{1})\Phi_{+}^{\dagger}(x_{1},x_{1})+\mathrm{e}^{\mathrm{i}kx_{1}}\xi^{-}_{j}(x_{1})\Phi_{-}^{\dagger}(x_{1},x_{1})\right)|\Omega\rangle$ (31)
and
$F(t)\equiv\langle\Phi_{2}|\mathrm{e}^{-\mathrm{i}H_{\textrm{iCFT}}t}|\Phi_{1}\rangle=\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\langle\Omega|\left(\mathrm{e}^{\mathrm{i}kx_{2}}\overline{\xi^{+}_{2}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(x_{2},x_{2})+\mathrm{e}^{-\mathrm{i}kx_{2}}\overline{\xi^{-}_{2}(x_{2})}\Phi_{-}^{\vphantom{\dagger}}(x_{2},x_{2})\right)\mathrm{e}^{-\mathrm{i}H_{\textrm{iCFT}}t}\\ \times\left(\mathrm{e}^{-\mathrm{i}kx_{1}}\xi^{+}_{1}(x_{1})\Phi_{+}^{\dagger}(x_{1},x_{1})+\mathrm{e}^{\mathrm{i}kx_{1}}\xi^{-}_{1}(x_{1})\Phi_{-}^{\dagger}(x_{1},x_{1})\right)|\Omega\rangle,$ (32)
in order to eventually compare $|F(t)|$ with $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for specular waves, $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$. Using the unfolding relations (23), we obtain
$\langle\Phi_{j}|\Phi_{j}\rangle=\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\langle\Omega|\left(\mathrm{e}^{\mathrm{i}kx_{2}}\overline{\xi^{+}_{j}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(x_{2},x_{2})+\mathrm{e}^{-\mathrm{i}kx_{2}}\overline{\xi^{-}_{j}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(L-x_{2},L-x_{2})\right)\\ \times\left(\mathrm{e}^{-\mathrm{i}kx_{1}}\xi^{+}_{j}(x_{1})\Phi_{+}^{\dagger}(x_{1},x_{1})+\mathrm{e}^{\mathrm{i}kx_{1}}\xi^{-}_{j}(x_{1})\Phi_{+}^{\dagger}(L-x_{1},L-x_{1})\right)|\Omega\rangle$ (33)
and
$F(t)=\mathrm{e}^{-\mathrm{i}Ct}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\langle\Omega|\mathrm{e}^{\mathrm{i}H_{\textrm{iCFT}}t}\left(\mathrm{e}^{\mathrm{i}kx_{2}}\overline{\xi^{+}_{2}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(x_{2},x_{2})+\mathrm{e}^{-\mathrm{i}kx_{2}}\overline{\xi^{-}_{2}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(L-x_{2},L-x_{2})\right)\mathrm{e}^{-\mathrm{i}H_{\textrm{iCFT}}t}\\ \times\left(\mathrm{e}^{-\mathrm{i}kx_{1}}\xi^{+}_{1}(x_{1})\Phi_{+}^{\dagger}(x_{1},x_{1})+\mathrm{e}^{\mathrm{i}kx_{1}}\xi^{-}_{1}(x_{1})\Phi_{+}^{\dagger}(L-x_{1},L-x_{1})\right)|\Omega\rangle,$ (34)
where we used that $H_{\textrm{iCFT}}|\Omega\rangle=C|\Omega\rangle$ for some real constant $C$ that encodes the Casimir energy. (This only gives rise to an overall phase that one could safely ignore.) Clearly, as stated in the main text, the above amounts to evaluating two-point correlations for primary fields, specifically those of right movers in a chiral CFT on $[-L/2,3L/2]$.
First, as a preliminary step, for primary fields with periodic boundary conditions on the system with length $2L$, conformal invariance implies that the undeformed two-point correlations are
$\langle\Omega_{0}|\Phi^{\vphantom{\dagger}}_{+}(x_{2}^{-},x_{2}^{+})\Phi^{\dagger}_{+}(x_{1}^{-},x_{1}^{+})|\Omega_{0}\rangle=\frac{1}{2\pi}G^{+}_{\Phi\Phi}(x_{2}^{-}-x_{1}^{-})^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}(x_{2}^{+}-x_{1}^{+})^{2\Delta^{-}_{\Phi}},$ (35)
where $G^{\pm}_{\Phi\Phi}(x)$ are $2L$-periodic functions. For example, as can be seen from, e.g., [55], the correlations for right-moving free fermions (cf. Example 1 above) on the system with length $2L$ and periodic BCs are
$\langle\Omega_{0}|\psi^{\vphantom{\dagger}}_{+}(x_{2})\psi^{\dagger}_{+}(x_{1})|\Omega_{0}\rangle=\frac{1}{2\pi}G^{+}_{\textrm{FF}}(x_{2}-x_{1})$ (36)
with
$G^{+}_{\textrm{FF}}(x)\equiv\frac{\mathrm{i}\pi\exp\bigl{(}-\mathrm{i}\frac{\pi}{2L}x\bigr{)}}{2L\sin\bigl{(}\frac{\pi}{2L}[x+\mathrm{i}0^{+}]\bigr{)}},$ (37)
where we used that $\Delta^{+}_{\psi}=1/2$, $\Delta^{-}_{\psi}=0$. Similarly, for the Luttinger model (cf. Example 2 above), the correlations for the (renormalized) right-moving fermionic fields are
$\langle\Omega_{0}|\psi^{\vphantom{\dagger}}_{+}(x_{2}^{-},x_{2}^{+})\psi^{\dagger}_{+}(x_{1}^{-},x_{1}^{+})|\Omega_{0}\rangle=\frac{1}{2\pi}G^{+}_{\textrm{FF}}(x_{2}^{-}-x_{1}^{-})^{(1+K)^{2}/4K}G^{+}_{\textrm{FF}}(x_{1}^{+}-x_{2}^{+})^{(1-K)^{2}/4K},$ (38)
where we used that
$G^{-}_{\textrm{FF}}(x)\equiv\frac{-\mathrm{i}\pi\exp\bigl{(}+\mathrm{i}\frac{\pi}{2L}x\bigr{)}}{2L\sin\bigl{(}\frac{\pi}{2L}[x-\mathrm{i}0^{+}]\bigr{)}}=\frac{\mathrm{i}\pi\exp\bigl{(}-\mathrm{i}\frac{\pi}{2L}[-x]\bigr{)}}{2L\sin\bigl{(}\frac{\pi}{2L}[-x+\mathrm{i}0^{+}]\bigr{)}}=G^{+}_{\textrm{FF}}(-x).$ (39)
Second, by results spelled out in Proposition 3.1 in [26] for inhomogeneous CFT, the time dependence in $F(t)$ means that those correlations are evaluated at spacetime points given by $\underline{x}^{\pm}_{t}(x)$ in Eq. (27). In addition, the action of $U_{+}(\underline{f})$ from the expression (29) for $|\Omega\rangle$ also leads to a conformal transformation of the fields. It follows from this and the properties (27) that
$\langle\Phi_{j}|\Phi_{j}\rangle=\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl{(}&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{j}(x_{2})}\xi^{+}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{j}(x_{2})}\xi^{-}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{j}(x_{2})}\xi^{+}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{j}(x_{2})}\xi^{-}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})\bigr{)}^{2\Delta^{-}_{\Phi}}\biggr{)}\end{aligned}$ (40)
and
$F(t)=\mathrm{e}^{-\mathrm{i}Ct}\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl{(}&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\biggr{)},\end{aligned}$ (41)
where we used that $\underline{f}\bigl{(}\underline{x}_{t}^{\pm}(x)\bigr{)}=\underline{f}(x)\pm\underline{v}_{0}t$ and $\underline{f}\bigl{(}\underline{x}_{t}^{\pm}(L-x)\bigr{)}=\underline{f}(L)-\underline{f}\bigl{(}\underline{x}_{t}^{\mp}(x)\bigr{)}=\underline{f}(L)-\underline{f}(x)\pm\underline{v}_{0}t$.
Finally, for specular waves, $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$, we obtain
$F(t)=\mathrm{e}^{-\mathrm{i}Ct}\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(-x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(-x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl{(}&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{1}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(-x_{2})-\underline{f}(x_{1})+\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(-x_{2})-\underline{f}(x_{1})+\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{1}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(-x_{2})+\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(-x_{2})+\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{1}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(-x_{2})-\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(-x_{2})-\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{1}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(-x_{2})+\underline{f}(x_{1})-\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(-x_{2})+\underline{f}(x_{1})-\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\biggr{)},\end{aligned}$ (42)
after a change of variable and rearranging.
By comparing the terms appearing in Eqs. (42) and (40), it follows that $|F(t)|$ for certain $t$ equals $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for all $\xi^{\pm}_{1}(x)$ if and only if $\underline{v}(x)$ is even, meaning $\underline{f}(x)$ must be odd. This in turn implies $\underline{f}(L)=L$ and $\underline{v}_{0}=v_{0}$ given by Eq. (7), which means that equality holds for all $t=(2m+1)L/v_{0}$ due the $2L$ periodicity of $G^{\pm}_{\Phi\Phi}(x)$. [Moreover, for $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$, one has $|\langle\Phi_{1}|\Phi_{1}\rangle|=|\langle\Phi_{2}|\Phi_{2}\rangle|$ when $v(x)$ is even.] Since $\underline{v}(x)=v(x)$ for all $x\in[-L/2,L/2]$, this proves the claim that one-particle PWT is exhibited for all even inhomogeneous CFTs.
As a last remark, one can unfold in a different way to take the spin/statistics of the primary fields into account, so that, e.g., one gets anti-periodic BCs for fermions on the system of size $2L$, and twisted BCs for general anyonic fields. This can be used to obtain simple explicit formulas for the two-point correlations of any primary fields. However, since we are interested in recombinations of waves traveling in different directions, it is necessary to identify so as to always have periodic BCs. In any case, for our purposes, all we needed was the spacetime dependence of the correlations, which are fixed on general grounds in CFT, circumventing any need for explicit formulas.
Part B: Spectral properties of inhomogeneous TLLs
In this part, we first show that reflection symmetry is a necessary condition for PWT for any theory; here under the assumption that the Hamiltonian has pure point spectrum. Second, we give the details for how to use SL theory to diagonalize the inhomogeneous TLL Hamiltonian (8) and to formulate the PWT requirements on the spectrum. Third, we review some details on Weyl’s law applied to SL operators and, fourth, extend the previous discussion to TLL Hamiltonians with a mass term. Fifth, we comment on the Liouville normal form used to state the results for the inverse SL problem in the main text.
Necessity of reflection symmetry
Before considering TLL theory, we note that all models studied in the main text have pure point spectrum [as long as the integral defining $1/v_{0}$ in Eq. (7) converges]. Thus, we first consider a general Hamiltonian $H$ with pure point spectrum and show that for such a system to exhibit PST/PWT, it is necessary that $H$ commutes with the parity operator $\mathcal{P}$, which maps $x\mapsto-x$. In other words, we will show that reflection symmetry is a necessary condition.
To prove the statement, let $E_{n}$ denote the eigenenergies of $H$ with the corresponding energy eigenstates $|n\rangle$. On general grounds, the latter form a complete orthonormal basis for the Hilbert space $\mathcal{F}$ of the theory. Let $|\Phi_{1}\rangle$ be an arbitrary state in $\mathcal{F}$ and let $|\Phi_{2}\rangle\equiv\mathcal{P}|\Phi_{1}\rangle$. (We recall that $\mathcal{P}$ is a unitary operator satisfying $\mathcal{P}^{2}=1$, meaning that $\mathcal{P}^{-1}=\mathcal{P}^{\dagger}=\mathcal{P}$.) As usual, define $F(t)\equiv\langle\Phi_{2}|\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle=\langle\Phi_{1}|\mathcal{P}\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle$. For general PST/PWT, we require that there exists $T>0$ such that $|F(T)|=|\langle\Phi_{1}|\Phi_{1}\rangle|$ for any state $|\Phi_{1}\rangle$. In particular, this must hold for any energy eigenstate $|n\rangle$, implying $|\langle n|\mathcal{P}|n\rangle|=1$. If $|n\rangle|$ is not an eigenstate of $\mathcal{P}$, it is straightforward to show that this equality cannot hold. Consequently, all energy eigenstates must also be eigenstates of $\mathcal{P}$, which implies that $[\mathcal{P},H]=0$.
Inhomogeneous TLLs and SL theory
We now turn to inhomogeneous TLLs studied in the main text. Inserting the expansions (14) into Eq. (10) yields
$H_{\textrm{iTLL}}=\frac{1}{2\pi}[\pi\Pi_{0}]^{2}+\frac{1}{2\pi}\sum_{n=1}^{\infty}\left.:\!\left([\pi\Pi_{n}]^{2}+\lambda_{n}\varphi_{n}^{2}\right)^{2}\!:\right.,$ (43)
where $\varphi_{0}$ is absent (when $E_{0}=0$), no different from the standard case with constant $K(x)$ [and $v(x)$]. For $n\geq 1$, let
$\varphi_{n}=\sqrt{\frac{\pi}{2}}\frac{1}{\lambda_{n}^{1/4}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right),\qquad\Pi_{n}=-\mathrm{i}\frac{\lambda_{n}^{1/4}}{\sqrt{2\pi}}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ (44)
in terms of bosonic creation and annihilation operators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$, satisfying $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$, as in the main text. Inserting Eq. (44) into Eq. (43) yields Eqs. (15)–(16).
Using the above, one can compute the time evolution of any observable in the algebra of operators generated by the fields $\varphi(x)$ and $\Pi(x)$. For instance, for $\varphi(x,t)\equiv\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}\varphi(x)\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}$ and $\Pi(x,t)$ defined similarly, we obtain
$\varphi(x,t)=\sum_{n=1}^{\infty}\sqrt{\frac{\pi}{2E_{n}}}\left(a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}+a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}\right)u_{n}(x),\qquad\Pi(x,t)=\sum_{n=1}^{\infty}(-\mathrm{i})\sqrt{\frac{E_{n}}{2\pi}}\left(a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}-a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}\right)w(x)u_{n}(x),$ (45)
where we omitted the zero modes (as they have to be treated separately when $E_{0}=0$) and used $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}$ and $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\dagger}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}$. Generally, any observable $\mathcal{O}(x,t)$ in this algebra will contain contributions of the form $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$. Now, in order for PWT to be exhibited, Eq. (1) must hold for any state, which means that it holds true without expectation values, i.e., $\mathcal{O}(x,0)$ must equal $\mathcal{O}(-x,T)$ at some later time $T>0$ as an operator identity. Taking $\mathcal{O}=\varphi$ and using the linear independence of the generators $a^{\vphantom{\dagger}}_{n}$ and $a^{\dagger}_{n}$, Eq. (45) yields
$u_{n}(x)=u_{n}(-x)\mathrm{e}^{\pm\mathrm{i}E_{n}T}\quad\forall n=0,1,\ldots$ (46)
for a common $T>0$, where we included the trivial case $n=0$ [recall that $u_{0}(x)=\textrm{const}$ and $E_{0}=0$ when $q(x)=0$ in Eq. (11)]. Since $\mathcal{P}u_{n}(x)=u_{n}(-x)$, Eq. (46) implies that each $u_{n}(x)$ must be an eigenfunction of the parity operator $\mathcal{P}$, whose eigenvalues are $+1$ or $-1$ (since $\mathcal{P}^{2}=1$). It follows that all eigenfunctions must have definite parity, and the only possibility for PWT to hold is that the corresponding eigenenergies satisfy $E_{n}T=\pi m_{n}$ for integers $m_{n}$ for each $n$, with $m_{n}$ even (odd) if $u_{n}(x)$ has even (odd) parity.
Finally, it is straightforward to show the following statement: The eigenfunctions $u_{n}(x)$ of a regular SL operator $\mathcal{A}$ have definite parity if and only if $[\mathcal{A},\mathcal{P}]=0$. For our operator in Eqs. (11)–(12), this means that both $v(x)$ and $K(x)$ must be even functions (using that they must be positive to exclude the possibility that both are instead odd).
Given the above, together with Properties (i)(iv) for regular SL operators stated in the main text, we have derived our main result on the spectral requirements for an inhomogeneous TLL to exhibit PWT.
Moreover, as stated in the main text, reversing the logic, namely, given a spectrum that satisfy the spectral properties for PWT to be exhibited, our formulation of PWT for inhomogeneous TLLs as an inverse SL problem follows.
Weyl’s law for SL operators
Weyl’s law applied to SL operators, see, e.g., [78, 79], states that the number $N(\lambda)$ of eigenvalues below a given $\lambda>0$ asymptotically behaves as
$\frac{N(\lambda)}{\sqrt{\lambda}}\sim\frac{1}{\pi}\int_{-L/2}^{L/2}\mathrm{d}x\,\sqrt{\frac{w(x)}{p(x)}}$ (47)
for large $\lambda$. This holds under certain technical conditions on the coefficients $w(x)$, $p(x)$ in Eq. (11) [which are satisfied assuming our $v(x)$, $K(x)$ are positive and smooth functions]. Using Eqs. (12) and (16) as well as Property (i) in the main text, it follows that
$\frac{n}{E_{n}}\sim\frac{1}{\pi}\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{v(x)}=\frac{L}{\pi v_{0}}$ (48)
for large $n$, where in the last step we used Eq. (7). Thus, assuming that the integral defining $1/v_{0}$ converges, then this gives the stated asymptotic behavior, $E_{n}=\pi v_{0}n/L+o(n)$, in the main text. (Another possibility would be if the integral diverges, in which case $E_{n}\sim\textrm{const}\times n^{\gamma}$ for $\gamma<1$, but this is inconsistent with that the eigenenergies must be proportional to integers $m_{n}$ with a common proportionality constant in order for the theory to exhibit PWT.)
“Massive” inhomogeneous TLLs
For massless TLL bosons, as discussed above and in the main text, given a spectrum satisfying the PWT requirements, we saw that $K(x)$ can be found uniquely up to rescaling and changing coordinates.
We now extend to SL operators (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq. (8):
$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\biggl{(}\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}+v_{0}^{2}M(x)^{2}\varphi(x)^{2}\biggr{)}\!:\right.\!,$ (49)
Indeed, it is straightforward to extend the previous discussion to such models. For instance, the PWT requirements on the spectrum are the same, except the spectrum does not start at zero: For $q(x)\neq 0$, a constant $u_{0}(x)$ no longer solves Eq. (13), thus $E_{0}>0$. Instead, we have a “massive” inhomogeneous TLL, for which one can show:
A “massive” inhomogeneous TLL exhibits PWT if and only if $v(x)$, $K(x)$, $q(x)$ are even
and all eigenvalues are of the form $E_{n}=\pi\tilde{m}_{n}/T$ for certain integers $\tilde{m}_{n}$.
Here, these integers $\tilde{m}_{n}=m_{n}+c$ in terms of the integers $m_{n}$ for the massless case in the main text, where we allow for an overall shift by a finite even integer $c$. [The latter must be even so that $\tilde{m}_{n}$ is even (odd) if $n$ is even (odd), as required due to Property (ii) in the main text.] Moreover, given such a spectrum, the results in [82, 83] still imply that $\hat{q}(y)$ is uniquely determined by the spectrum, but this can no longer uniquely determine $K(x)$ since the first term in Eq. (18) is now non-zero in general.
As an example, we revisit the case with $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl{[}1-(2x/L)^{2}\bigr{]}^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$ encountered in the main text, whose eigenfunctions were given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$, but now we include a fine-tuned potential $q(x)=-(2v\alpha/L)^{2}w(x)$, where we recall that $w(x)=K(x)/v(x)$. This fine-tuned potential allows us to complete the square so that the eigenvalues in Eq. (13) are $\lambda_{n}=(\pi v_{0}/L)(n+\alpha)^{2}$ with $v_{0}=2v/\pi$, meaning that $E_{n}=\pi(n+\alpha)v_{0}/L$. Consequently, any $c=\alpha=0,2,4,\ldots$ works, yielding a discrete family of models with fine-tuned inhomogeneous mass that exhibit PWT. We stress that the mass in this example satisfies $M(x)\propto 1/L$ and therefore vanishes in the thermodynamic limit. Thus, one should not expect hallmarks of massive theories, such as exponential decay of two-point correlations with respect to the spatial separation.
Liouville normal form and compactified free bosons
It is curious to note the following “shortcut” to writing our massless bosons (8) in Liouville normal form: Recall that an inhomogeneous TLL is a theory of compactified free bosons with a position-dependent compactification radius $R(x)$ in curved spacetime given by the metric $(h_{\mu\nu})=\operatorname{diag}(v(x)^{2}/v_{0}^{2},-1)$ for coordinates $(x^{0},x^{1})\equiv(v_{0}t,x)$; see, e.g., [27, 53]. Given a Luttinger parameter profile $K(x)$, the compactification radius is $R(x)=\sqrt{2\alpha^{\prime}}K(x)$ ($\alpha^{\prime}$ has dimension length squared and is commonly set equal to $2$). In terms of the bosonic field $\varphi(x)$ (with values taken modulo $2\pi$), the action corresponding to the Hamiltonian (8) is
$S=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}R(x)^{2}(\partial_{\mu}\varphi)(\partial^{\mu}\varphi),$ (50)
where we recall that $h\equiv\det(h_{\mu\nu})$. This can be viewed as a non-exactly marginal deformation of a theory with constant compactification radius; cf. [53]. The corresponding action for $X(x)\equiv R(x)\varphi(x)$ becomes
$S=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}\bigl{[}(\partial_{\mu}X)(\partial^{\mu}X)-\hat{q}X^{2}/v(x)^{2}\bigr{]}$ (51)
with $\hat{q}=\hat{q}(y)$ in Eq. (18) for $y=y(x)$ given by Eq. (7), which agrees with Eq. (19) for $\mathfrak{u}(y)=X(x)/\sqrt{2\alpha^{\prime}}$. I.e., as a theory for $X(x)$, even when $q(x)=0$, the position-dependent $K(x)$ corresponds a (possibly imaginary) inhomogeneous mass $\sqrt{\hat{q}(y)}/v(x)$. This also extends to the case
$S=\frac{1}{2\pi}\int\mathrm{d}^{2}x\sqrt{-h}\bigl{[}K(x)(\partial_{\mu}\varphi)(\partial^{\mu}\varphi)-v_{0}^{2}M(x)^{2}\varphi^{2}\bigr{]}$ (52)
with a non-zero mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ for the bosonic field $\varphi(x)$, corresponding to the Hamiltonian (49), the only difference being that the first term in Eq. (18) is then no longer zero.
Part C: Correlation functions in inhomogeneous TLLs
Consider the inhomogeneous TLL theory with Hamiltonian $H_{\textrm{iTLL}}$ in Eq. (8). Using the eigenfunctions obtained from SL theory, we can compute any ground-state two-point correlation functions for the fields $\varphi(x)$ and $\vartheta(x)$, where the latter is defined by
$\partial_{x}\vartheta(x)\equiv\pi\Pi(x).$ (53)
Indeed, using the expansions (45), that $a_{n}$ annihilates the ground state $|\Omega\rangle$ of $H_{\textrm{iTLL}}$, as well as the commutation relations $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$, it is straightforward to show that
$\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ $=\sum_{n=1}^{\infty}\frac{\pi}{2E_{n}}u_{n}(x)u_{n}(x^{\prime})\mathrm{e}^{-\mathrm{i}E_{n}(t-t^{\prime})},$ (54)
$\langle\Omega|\varphi(x,t)\vartheta(x^{\prime},t^{\prime})|\Omega\rangle$ $=\mathrm{i}\sum_{n=1}^{\infty}\frac{\pi}{2}u_{n}(x)U_{n}(x^{\prime})\mathrm{e}^{-\mathrm{i}E_{n}(t-t^{\prime})},$
$\langle\Omega|\vartheta(x,t)\vartheta(x^{\prime},t^{\prime})|\Omega\rangle$ $=\sum_{n=1}^{\infty}\frac{\pi E_{n}}{2}U_{n}(x)U_{n}(x^{\prime})\mathrm{e}^{-\mathrm{i}E_{n}(t-t^{\prime})},$
where we introduced $U_{n}(x)$ defined by
$\partial_{x}U_{n}(x)\equiv w(x)u_{n}(x).$ (55)
Similar correlation functions can be computed for other objects, including vertex operators (i.e., normal-ordered exponentials of the bosonic fields) [54, 55, 56, 57, 58, 59]. It is also straightforward to compute Green’s functions; see, e.g., [27], the only difference with the present case being that one must account for the non-trivial weight function $w(x)$ that was simply a constant in [27].
As an important example, consider a theory with $K(x)=K>0$ constant but $v(x)>0$ general [and $q(x)=0$]. We recall that the SL operator in this case is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq. (7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl{(}n\pi[L+2y]/2L\bigr{)}$ and $u_{0}(x)=\sqrt{v_{0}/L}$. Then, one can show that
$\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle\\ \begin{aligned} =\frac{1}{4}&\left(\log\frac{\mathrm{i}\exp\Bigl{(}-\mathrm{i}\frac{\pi}{2L}[f(x)-f(x^{\prime})-v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)-f(x^{\prime})-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr{)}}+\log\frac{\mathrm{i}\exp\Bigl{(}-\mathrm{i}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L-v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr{)}}\right.\\ &\quad+\left.\log\frac{-\mathrm{i}\exp\Bigl{(}\mathrm{i}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L+v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr{)}}+\log\frac{-\mathrm{i}\exp\Bigl{(}\mathrm{i}\frac{\pi}{2L}[f(x)-f(x^{\prime})+v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)-f(x^{\prime})+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr{)}}\right),\end{aligned}$ (56)
where we included the necessary regularizations in order to interpret coincident spacetime points and used the identity $\sum_{n=1}^{\infty}\frac{1}{n}\mathrm{e}^{-\xi n}=-\log\left(1-\mathrm{e}^{-\xi}\right)$ for any $\xi$ such that $\operatorname{Re}\xi>0$. Eq. (56) is the analytical expression for what was plotted in Fig. 1 (assuming $K=1$). Moreover, for delta-like waves, the terms appearing in $\log F(t)$ given by Eq. (41) for the free-fermion case (37) closely resemble those in Eq. (56). Indeed, the former agrees with $4\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ to leading order, and one can furthermore show that the agreement between the two ways of computing (CFT methods versus bosonization) is exact (as it should be) by evaluating and summing all combinations of two-point correlation functions of the vertex operators describing the right/left-moving free fermionic fields.

### v3

@@
 Abstract
-We study the perfect transfer of information in 1+1D continuous quantum systems.
-This includes effective descriptions of inhomogeneous spin chains, for which the notion of perfect state transfer in quantum information was introduced, and here phrased in terms of waves.
-We show that reflection symmetry is necessary for perfect wave transfer (PWT) in any inhomogeneous conformal field theory, and even sufficient when restricted to one-particle excitations.
-To determine if or when it is sufficient more generally, we first break conformal invariance and study a broad class of 1+1D bosonic theories.
-We show that the question can then be posed as an inverse Sturm-Liouville problem that determines when the bosonic theory exhibits PWT.
-We demonstrate how to uniquely solve this problem, which also shows that reflection symmetry is sufficient for the special case with conformal invariance.
-Using bosonization, our continuum results extend these notions to interacting quantum field theories.
-Introduction.—Non-photonic quantum communication enables the efficient transmission of information from one quantum register to another [1].
-This is desirable for quantum networks and computation [2, 3], not least since it allows for the perfect transfer of states even without dynamical control [4, 5].
-First shown theoretically for communication channels in the form of a spin chain with spatially varying hopping, such perfect state transfer (PST) was extended to spins on graphs and other networks [6, 7, 8, 9, 10, 11, 12, 13, 14] and demonstrated experimentally [15, 16, 17, 18, 19, 20].
-However, the notion has remained squarely in the discrete realm.
-In this paper, we extend PST to continuous systems in one space and one time dimension (1+1D).
-In so doing, a rephrasing in terms of waves becomes natural, motivating the term perfect wave transfer (PWT).
-As for states in PST, a system on the interval $[-L/2,L/2]$ of length $L$ with suitable boundary conditions (BCs) exhibits PWT if the propagation of any initial wave, after some time $T$, amounts to its reflection; see Fig.
+The transfer of information from one part of a quantum system to another is fundamental to the understanding and design of quantum information processing devices.
+In the realm of discrete systems such as spin chains, inhomogeneous networks have been engineered that allow for the perfect transfer of qubits from one end to the other.
+Here, by contrast, we investigate the perfect transfer of information in continuous systems, phrased in terms of wave propagation.
+A remarkable difference is found between systems that possess conformal invariance and those that do not.
+Systems in the first class enjoy perfect wave transfer (PWT), explicitly shown for one-particle excitations and anticipated in general.
+In the second class, those that exhibit PWT are characterized as solutions to an inverse spectral problem.
+As a concrete example, we demonstrate how to formulate and solve this problem for a prototypical class of bosonic theories, showing the importance of conformal invariance for these theories to enjoy PWT.
+Using bosonization, our continuum results extend to theories with interactions, broadening the scope of perfect information transfer to more general quantum systems.
+Introduction—A fundamental task in quantum information processing is to transfer information from one register to another [1, 2].
+Optical fibers may be used, but efficiency can be improved using a quantum wire as both communication channel and registers [3].
+Since this relies on collective phenomena to transfer quantum states, efficiency is optimized by engineering the system’s many-body properties.
+This allows not only for high-fidelity but even perfect transfer of states without dynamical control of individual sites.
+Such perfect state transfer (PST) was first shown theoretically for wires modeled as spin chains with inhomogeneous hopping [4, 5], then extended to spins on graphs and other networks [6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and demonstrated experimentally [16, 17, 18, 19, 20, 21].
+However, PST has so far been addressed only in the discrete realm, while powerful continuum frameworks to describe collective quantum phenomena remain untapped.
+In this Letter, we harness that power to investigate quantum information transfer in inhomogeneous continuous systems in 1+1 dimensions (1+1D).
+At low energies, such examples are realizable using, e.g., spin ladders [22], Josephson-junction arrays [23], ultracold atoms [24], and nanowires [25, 26, 27].
+In this framework, rephrasing using waves becomes natural, motivating the term perfect wave transfer (PWT): As for states in PST, a system on an interval [L/2,L/2][-L/2,L/2] of length $L$ with suitable boundary conditions (BCs) exhibits PWT if propagating any initial wave for some time $T$ amounts to a reflection; see Fig.
 1.
-Concretely: The system enjoys PWT if
-$\langle\mathcal{O}(x,0)\rangle=\langle\mathcal{O}(-x,T)\rangle,$ (1)
-[…]
-Related recent advances in the continuum include non-equilibrium studies of inhomogeneous versions of conformal field theory (CFT).
-These are conformally-invariant quantum field theories deformed so that the usual constant propagation velocity becomes a profile $v(x)$ [21, 22, 23, 24, 25, 26].
-Among their applications, such inhomogeneous CFTs effectively describe gapless quantum many-body systems with mesoscopic inhomogeneities, such as spin chains with spatially varying hopping, trapped ultracold atoms, and arrays of quantum circuits; cf.
-[27].
-Special cases include the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$ for some constant $v>0$ [28, 29]—the continuum counterpart of the inhomogeneous spin chain in [4] that first exhibited PST—and profiles produced by Möbius transformations [30, 31, 32, 33, 34, 35, 36, 37].
-Note that for PWT, as for other transport properties of finite systems, the BCs [38, 39, 29, 40] are crucial.
-Furthermore, inhomogeneous CFT can be viewed as a theory in curved spacetime.
-As such, besides applications to quantum communication in solid-state devices [41, 42], our work also relates to relativistic quantum information [43, 44].
-Since the same theories are also natural to study on the half or full infinite line, our results may likewise be relevant for long-distance quantum communication [45].
 Figure 1: (a) Illustration of PWT with Neumann BCs showing an initial wave that evolves to its reflection at a certain time $T$.
@@
 (7).
-Specifically, the magnitude of boson two-point correlations (17) at the free-fermion point is plotted, which by bosonization encodes the same information as $\log F(t)$ in Eq.
-(5) for delta-like waves $\xi_{1}$ centered at $-3L/8$ and $\xi_{2}$ at $x$ (setting $k_{\textrm{F}}=0$ for simplicity and $t^{\prime}=0$).
-Engineering PST/PWT is at heart a spectral problem, amounting to whether the propagator coincides with the parity operator at certain times.
-We address this engineering question for quantum field theories describing a continuous quantum communication channel.
-Remarkably, for bosonic theories [46, 47, 48, 49, 50, 51, 52, 53], we arrive at an inverse problem for Sturm-Liouville operators, which we solve and whose solution determines if the channel enjoys PWT.
-Using bosonization [54, 55, 56, 57, 58, 59], our results extend the notion of PST/PWT to a large class of systems, including interacting fermionic quantum field theories.
-Inhomogeneous CFT.—To begin with, we study the continuum limit of the inhomogeneous spin chains in [1, 4].
-In the gapless regime, the continuous counterpart is an inhomogeneous CFT with a velocity profile $v(x)>0$, $x\in[-L/2,L/2]$ [21, 22, 23, 24, 25, 26].
-The simplest example is that of inhomogeneous free fermions, given by the Hamiltonian (up to an overall constant and setting $\hbar=1$)
-$H_{\textrm{iFF}}\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\frac{v(x)}{2}\Bigl{[}\!\left.:\!\psi^{\dagger}_{+}(-\mathrm{i})\partial_{x}\psi^{\vphantom{\dagger}}_{+}\!:\right.+\left.:\!\psi^{\dagger}_{-}\mathrm{i}\partial_{x}\psi^{\vphantom{\dagger}}_{-}\!:\right.+\mathrm{h.c.}\Bigr{]}$ (2)
+Specifically, the real part of boson two-point correlations (17) at the free-fermion point is plotted, which by bosonization encodes the same information as $\log F(t)$ in Eq.
+(5) for deltalike waves $\xi_{1}$ centered at 3L/8-3L/8 and $\xi_{2}$ at $x$ (setting $k_{\textrm{F}}=0$ for simplicity and $t^{\prime}=0$).
+Concretely, there is PWT if
+$\langle\mathcal{O}(x,0)\rangle=\langle\mathcal{O}(-x,T)\rangle$ (1)
+for all time-evolved observables $\mathcal{O}(x,t)$ and expectation values $\langle\cdot\rangle$ with respect to any state.
+The dynamics will be time independent, but the problem is also interesting for driven systems [19, 20, 21], with randomness [28], or relaxed to a property for specific classes of observables or states.
+Ultimately, our motivation for understanding PWT is the efficient high-fidelity transfer of information—a problem that has attracted recent attention, e.g., in the design of distributed networks for quantum computation [29].
+Related advances include nonequilibrium studies of inhomogeneous versions of conformal field theory (CFT), i.e., conformally invariant quantum field theories [30] deformed so that the usual constant propagation velocity becomes a profile $v(x)$ [31, 32, 33, 34, 35, 36].
+Among their applications, such inhomogeneous CFTs effectively describe gapless quantum many-body systems with mesoscopic inhomogeneities, e.g., spin chains with spatially varying hopping and trapped ultracold atoms; cf.
+[37].
+Special cases include the profile $v(x)=v\sqrt{1-(2x/L)^{2}}$ for some constant $v>0$ [38, 39]—the continuum counterpart of the inhomogeneous spin chain [4] first exhibiting PST—and profiles produced by Möbius transformations [40, 41, 42, 43, 44, 45, 46, 47].
+Note that for PWT, as for other transport properties in finite systems, the choice of BCs [48, 49, 39, 50] is important.
+Engineering PWT will, at heart, be a spectral problem, amounting to whether the propagator coincides with the parity operator at certain times.
+This is shown for prototypical quantum field theories describing an inhomogeneous quantum wire—our continuous channel—focusing on universal low-energy descriptions where conformal invariance can be parametrically turned on or off.
+Notably, these are directly applicable to ultracold atoms [51, 52, 49], while such experimental systems are even fundamentally continuous [53], in itself warranting a continuum treatment of information transfer.
+Remarkably, for bosonic theories [54, 55, 56, 57, 58, 59, 60, 61], we arrive at an inverse spectral problem for Sturm-Liouville operators.
+Its solution determines if the channel enjoys PWT, and we show that conformal invariance is important if the system is required to be sufficiently regular.
+Using bosonization [62, 63, 64, 65, 66, 67], our results extend PWT or PST to a large class of theories, including interacting fermions, and we also discuss a proposal to extend beyond low energies.
+Lastly, in addition to quantum communication in solid-state devices [68, 69], our PWT results relate to relativistic quantum information [70, 71], and potentially to communication over long distances [72].
+Inhomogeneous CFT—To begin, we study the continuum limit of inhomogeneous spin chains [3, 4].
+In the gapless regime, their continuous counterpart is an inhomogeneous CFT with a velocity profile $v(x)>0$ [31, 32, 33, 34, 35, 36].
+The simplest example is that of inhomogeneous free fermions, given by the Hamiltonian (setting $\hbar=1$) [73]
+$H_{\textrm{iFF}}\equiv\int_{-L/2}^{L/2}\hskip-2.3pt\mathrm{d}x\,\frac{v(x)}{2}\Bigl[\!\left.:\!\hskip-0.5pt\psi^{\dagger}_{+}(-\mathrm{i})\partial_{x}\psi^{\vphantom{\dagger}}_{+}\hskip-0.5pt\!:\right.+\left.:\!\hskip-0.5pt\psi^{\dagger}_{-}\mathrm{i}\partial_{x}\psi^{\vphantom{\dagger}}_{-}\hskip-0.5pt\!:\right.+\mathrm{h.c.}\Bigr]$ (2)
 […]
-We assume open (Neumann) BCs at $x=\pm L/2$.
-These can be incorporated by standard unfolding methods; cf.
-[61, 62, 57, 58, 38, 63]: Couple right and left movers by identifying $\psi_{-}(x)=\psi_{+}(L-x)$, giving a chiral theory of only right-moving fermions on $[-L/2,3L/2]$ with periodic BCs and velocity $v(x)$ for $x\in[-L/2,L/2]$ and $v(L-x)$ for $x\in[L/2,3L/2]$.
-This implies zero particle flow at $x=\pm L/2$ since the two fermion densities $\rho_{\pm}(x)=\left.:\!\psi^{\dagger}_{\pm}(x)\psi^{\vphantom{\dagger}}_{\pm}(x)\!:\right.$ are equal at each boundary and the current is $\propto\rho_{+}-\rho_{-}$.
-PST was probed in [4] using a time-dependent overlap between one-particle states that are reflections of each other. Its continuum counterpart is to study states created by fields
+PST was probed in [4] using a time-dependent overlap between one-particle states that were mutual reflections. Their continuum counterparts are states created by fields
 $\psi(x)\equiv\mathrm{e}^{\mathrm{i}k_{\textrm{F}}x}\psi_{+}(x)+\mathrm{e}^{-\mathrm{i}k_{\textrm{F}}x}\psi_{-}(x),$ (3)
-where $k_{\textrm{F}}$ denotes the Fermi momentum.
+where $k_{\textrm{F}}$ denotes the Fermi momentum [79]. As a subset of possible states, consider
+$|\Psi_{j}\rangle\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\xi_{j}(x)\psi^{\dagger}(x)|\Omega\rangle\qquad(j=1,2)$ (4)
+for smooth enough functions $\xi_{j}(x)$, interpreted as “waves” of fermionic excitations in each state, and study the time-dependent overlap
+$F(t)\equiv\langle\Psi_{2}|\mathrm{e}^{-\mathrm{i}H_{\textrm{iFF}}t}|\Psi_{1}\rangle.$ (5)
+The states in Eq. (4) form a reduced class, as only one-particle excitations are possible, and right and left movers are equally distributed. (The latter restriction is lifted in the SM [74].)
+To probe PWT for one-particle excitations as in [4], we restrict to specular waves, $\xi_{2}(x)=\xi_{1}(-x)$, and compare $F(t)$ in Eq.
+(5) with $\langle\Psi_{1}|\Psi_{1}\rangle$.
+Specifically, there is one-particle PWT if there exists $T>0$ so that
+$|F(T)|=|\langle\Psi_{1}|\Psi_{1}\rangle|$ (6)
+for all initial waves $\xi_{1}(x)$ in Eq. (4) [80].
 Eq.
-(3) corresponds to the low-energy linearization of the spin-chain dispersion relation around its Fermi points ±$\pm k_{\textrm{F}}$, followed by recentering the resulting fields so they describe right/left-moving excitations around zero momentum.
-As a preparatory subset of possible states, consider
-$|\Psi_{j}\rangle\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\xi_{j}(x)\psi^{\dagger}(x)|\Omega\rangle\qquad(j=1,2)$ (4)
-for smooth enough functions $\xi_{j}(x)$, interpreted as distributions or “wave packets” of fermionic excitations in each state, and study the time-dependent overlap
-$F(t)\equiv\langle\Psi_{2}|\mathrm{e}^{-\mathrm{i}H_{\textrm{iFF}}t}|\Psi_{1}\rangle.$ (5)
-The states (4) form a restricted class as only one-particle excitations are possible and as right/left movers have the same particle distributions; cf.
-the more general treatment in the SM [60] where we lift the latter restriction.
-To probe PWT for one-particle excitations as done in [4], we restrict to specular waves, $\xi_{2}(x)=\xi_{1}(-x)$, and compare $F(t)$ in Eq.
-(5) with $\langle\Psi_{1}|\Psi_{1}\rangle$.
-We say that: There is one-particle PWT if there exists $T>0$ so that
-$|F(T)|=|\langle\Psi_{1}|\Psi_{1}\rangle|$ (6)
-for all initial waves $\xi_{1}(x)$ in Eq. (4)  [^1] If Eq. (6) holds for any state $|\Psi_{1}\rangle$ in the Fock space, then general PST/PWT (i.e., for any number of excitations) would be enjoyed..
-Eq.
-(6) amounts to a comparison between two-point correlation functions at equal and non-equal times.
-Such can be computed in any inhomogeneous CFT by mapping them to the corresponding homogeneous correlations [26].
-The details are given in the SM [60] and boil down to comparing spacetime points after conformal transformations to coordinates given by
+(6) amounts to a comparison between two-point correlation functions at equal and nonequal times.
+Such correlations can be computed in any inhomogeneous CFT by mapping them to the corresponding homogeneous correlations [36].
+The details given in the SM [74] distill to comparing spacetime points after conformal transformations to coordinates
 $y=y(x)\equiv\int_{0}^{x}\mathrm{d}s\,\frac{v_{0}}{v(s)},\qquad\frac{1}{v_{0}}\equiv\frac{1}{L}\int_{-L/2}^{L/2}\frac{\mathrm{d}s}{v(s)}.$ (7)
-The outcome of the comparison shows that reflection symmetry is essential:
-An inhomogeneous CFT exhibits one-particle PWT
-if and only if $v(x)$ is even.
-The corresponding times are odd multiples, $(2m+1)T$ for any integer $m$, of the time $T=L/v_{0}$ that excitations take to cross the entire system’s length.
-For later usage, we note that $y\in[-L/2,L/2]$ if $v(x)$ is even.
-As an example, consider the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$, which gave the effective low-energy theory of the spin chain in [4].
-This continuum theory exhibits PWT, just like for any even profile $v(x)$, in contrast to [4] where that profile was exceptional in this respect.
+The comparison’s outcome shows that reflection symmetry is essential: Inhomogeneous CFTs exhibit one-particle PWT if and only if $v(x)$ is even.
+Moreover, this transfer occurs at odd multiples $(2m+1)T$ (for $m=0,1,\ldots$) of the time $T=L/v_{0}$ for excitations to cross the system’s entire length.
+Note that $y\in[-L/2,L/2]$ if $v(x)$ is even.
+As an example, consider the square-root profile $v(x)=v\sqrt{1-(2x/L)^{2}}$, effectively describing the spin chain in [4].
+It exhibits PWT, just like for any even profile $v(x)$, in contrast to [4], where that profile was exceptional in this respect.
 Notably, the square-root profile is positive everywhere except at $x=\pm L/2$.
 Thus, strictly speaking, $v(x)$ should be regularized before using the tools in [36], but since the integral defining $1/v_{0}$ in Eq.
-(7) converges, it has no consequence for PWT.
-Oppositely, the limiting case of inhomogeneous CFT given by Möbius transformations so that $v(x)$ decays to zero at the boundaries cannot exhibit PWT since $1/v_{0}$ diverges.
-Bosonic theories.—The above result and observations should not come as a surprise.
-After all, if one tries to compare discrete and continuous theories, then an inhomogeneous CFT can only describe low-energy excitations of a given lattice model.
-Moreover, one can show [60] that any Hamiltonian (with pure point spectrum) must commute with the parity operator for PWT to be possible, i.e., $v(x)$ must be even.
-The general structure of the spectrum due to conformal invariance in 1+1D [65] opens up the possibility that it might also be sufficient; e.g., the example (2) has a linear spectrum and is one-particle reducible, suggesting that if PWT holds for one-particle excitations then this should extend to many particles.
-The remainder of this paper concerns to what extent this is true more generally.
-Namely: What are necessary and sufficient conditions for PWT (1) for any number of excitations?
-To answer this question, we first break conformal invariance.
-Examples include 1+1D fermions with non-local interactions [66, 67, 68], while here we focus on bosons, specifically inhomogeneous Tomonaga-Luttinger liquids (TLLs) [46, 47, 48, 49, 50, 51, 52, 53].
-This theory is formulated using a compactified bosonic field $\varphi(x)$ (modulo $2\pi$) with conjugate $\Pi(x)$ satisfying $[\partial_{x}\varphi(x),\Pi(x^{\prime})]=\mathrm{i}\partial_{x}\delta(x-x^{\prime})$.
-The Hamiltonian is (up to an overall constant)
-$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\biggl{(}\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}\biggr{)}\!:\right.\!,$ (8)
-where, besides $v(x)>0$, the so-called Luttinger parameter $K(x)>0$ also depends on $x$.
+(7) converges, it has no consequence for PWT [81].
+Bosonic theories—The above result and observations should not come as a surprise.
+Indeed, if one compares discrete and continuous theories, then an inhomogeneous CFT can describe only low-energy excitations of a given lattice model.
+Moreover, we show [74] that any Hamiltonian (with pure point spectrum) must commute with the parity operator for PWT to be possible, i.e., $v(x)$ must be even, as should be expected.
+Furthermore, since excitations propagate ballistically in CFT, one might anticipate that this is not only necessary but also sufficient to ensure PWT with respect to all states.
+E.g., the model (2) has a linear spectrum and is free, suggesting that PWT for one-particle excitations extends to many particles.
+The rest of this Letter concerns to what extent this is true more generally: What are necessary and sufficient conditions for PWT as defined in Eq.
+(1)?
+To answer this question and understand the role of conformal invariance, we should not assume the latter.
+Specific nonconformal examples include 1+1D fermions with nonlocal interactions [82, 83, 84] or the Lieb-Liniger model [85].
+While interesting in their own right, we address this question using a paradigmatic low-energy theory, namely, inhomogeneous Tomonaga-Luttinger liquids (TLLs) [54, 55, 56, 57, 58, 59, 60, 61].
+By bosonization, this covers fermionic models generalizing the one in Eq.
+(2), while manipulating the inhomogeneities allows one to controllably turn conformal invariance on or off.
+The theory is formulated using a compactified bosonic field $\varphi(x)$ (modulo $2\pi$) with conjugate $\Pi(x)$ satisfying $\partial_{x}\varphi(x),\Pi(x^{\prime})]=\mathrm{i}\partial_{x}\delta(x-x^{\prime})$.
+The Hamiltonian is [73]
+$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\hskip-0.5pt\biggl(\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}\biggr)\hskip-0.5pt\!:\right$.\!, (8)
+where, in addition to $v(x)>0$, the Luttinger parameter $K(x)>0$ also depends on $x$.
 For simplicity, we assume that both functions are smooth.
-When $K(x)=K$ is constant, Eq.
-(8) is an inhomogeneous CFT, and when $K=1$, it is the bosonized version of the free fermions (2) [54, 55, 56, 57, 58, 59].
-However, $K^{\prime}(x)\neq 0$ locally couples right and left movers non-trivially, thus they do not commute and the theory is not conformal [53].
-Again, we impose open BCs, namely zero current at the boundaries:
-$\jmath(x)\big{|}_{x=\pm L/2}=0,\quad\textnormal{where}\;\;\jmath(x)\equiv-\frac{v(x)K(x)}{\pi}\partial_{x}\varphi(x)$ (9)
-is the total particle current. Unless stated otherwise, we also assume that $K^{\prime}(x)\big{|}_{x=\pm L/2}=0$ so that right and left movers are only coupled via the BCs (9) at those points.
-Sturm-Liouville (SL) theory.—Following [69, 70, 71, 72, 73, 74, 27], we use SL theory to find eigenfunctions, using which the fields can be expanded and the Hamiltonian (8) diagonalized.
-(This generalizes [27] as $v(x),K(x)$ are unrelated and differs from the approach in [53].) Indeed, using integration by parts and Eq.
+For $K(x)=K$ constant, Eq.
+(8) is an inhomogeneous CFT, and if $K=1$, it is the bosonized version of the free-fermion model (2) [62, 63, 64, 65, 66, 67].
+However, $K^{\prime}(x)\neq 0$ locally couples right and left movers nontrivially; thus, they do not commute, and the theory is not conformal [61].
+Again, we impose open BCs:
+$\hskip-2.84526pt\jmath(x)\big|_{x=\pm L/2}=0,\;\;\textnormal{where}\;\;\jmath(x)\equiv-\frac{v(x)K(x)}{\pi}\partial_{x}\varphi(x)$ (9)
+is the particle current. As a rule, we further assume that $K^{\prime}(x)$ is zero at $x=\pm L/2$ so that right and left movers are coupled only via the BCs (9) at those points [86].
+Sturm-Liouville (SL) theory—Following [87, 88, 89, 90, 91, 92, 37], we use SL theory to find eigenfunctions, using which the fields can be expanded and the Hamiltonian (8) diagonalized [93].
+Indeed, using integration by parts and Eq.
 (9),
-$H_{\textrm{iTLL}}=\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}w(x)\!\left.:\!\Biggl{(}\biggl{[}\frac{\pi\Pi(x)}{w(x)}\biggr{]}^{2}+\varphi(x)\mathcal{A}\varphi(x)\Biggr{)}\!:\right.\!$ (10)
+$H_{\textrm{iTLL}}=\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}w(x)\!\left.:\!\hskip-0.5pt\Biggl(\biggl[\frac{\pi\Pi(x)}{w(x)}\biggr]^{2}+\varphi(x)\mathcal{A}\varphi(x)\Biggr)\hskip-0.5pt\!:\right$.\!
+(10)
 with the SL operator
 $\mathcal{A}\equiv-\frac{1}{w(x)}\bigl[\partial_{x}p(x)\partial_{x}+q(x)\bigr],$ (11)
-$w(x)\equiv\frac{K(x)}{v(x)},\quad p(x)\equiv v(x)K(x),\quad q(x)\equiv 0,$ (12)
-where $\mathcal{A}$ acts on the space $\mathcal{D}$ of smooth  [^2] More precisely, twice continuously differentiable.
-functions $u(x)$ on $[-L/2,L/2]$ that satisfy $v(x)K(x)\partial_{x}u(x)\big{|}_{x=\pm L/2}=0$, i.e., Neumann BCs when $v(x),K(x)>0$.
+$w(x)\equiv\frac{K(x)}{v(x)},\quad p(x)\equiv v(x)K(x),\quad q(x)=0,$ (12)
+where $\mathcal{A}$ acts on the space $\mathcal{D}$ of smooth [94] functions $u(x)$ on [L/2,L/2][-L/2,L/2] that satisfy $v(x)K(x)\partial_{x}u(x)\big|_{x=\pm L/2}=0$, i.e., Neumann BCs when $v(x),K(x)>0$.
 We want to solve the SL problem
 $\mathcal{A}u=\lambda u,\quad u\in\mathcal{D}.$ (13)
-A distinction is warranted: Eq.
-(13) is regular if certain technical conditions are satisfied  [^3] Here, these amount to $v(x)$, $K(x)$ being strictly positive and $v(x)$, $K(x)$, $\partial_{x}p(x)$ [and $q(x)$ if non-zero] continuous on the whole interval $[-L/2,L/2]$.; otherwise it is irregular.
-Its importance stems from properties that hold for regular problems [77]:
+It is regular if certain conditions are satisfied [95], and irregular otherwise. This matters due to properties that hold for regular problems [96]:
   1. (i)
-    The spectrum consists of eigenvalues $\lambda_{n}$ that are real and ordered: $0\leq\lambda_{0}<\lambda_{1}<\ldots$.
+    The spectrum consists of eigenvalues $\lambda_{n}$ that are real and ordered so that $0\leq\lambda_{0}<\lambda_{1}<\ldots$.
   2. (ii)
-    Each $\lambda_{n}$ has a unique square-integrable eigenfunction $u_{n}(x)$ with $n$ isolated zeros in $[-L/2,L/2]$.
+    Each $\lambda_{n}$ has a unique square-integrable eigenfunction $u_{n}(x)$ with $n$ isolated zeros in [L/2,L/2][-L/2,L/2].
   3. (iii)
-    The $u_{n}(x)$ form a complete orthonormal basis for square-integrable functions on $[-L/2,L/2]$ with weight $w(x)$: $\int_{-L/2}^{L/2}\mathrm{d}x\,w(x)u_{n}(x)u_{n^{\prime}}(x)=\delta_{n,n^{\prime}}$.
-Furthermore, assuming that the integral defining $1/v_{0}$ in Eq. (7) converges:
+    The $u_{n}(x)$ form a complete orthonormal basis for square-integrable functions on [L/2,L/2][-L/2,L/2] with weight $w(x)$, i.e., $\int_{-L/2}^{L/2}\mathrm{d}x\,w(x)u_{n}(x)u_{n^{\prime}}(x)=\delta_{n,n^{\prime}}$.
+Furthermore, assuming that the integral defining $1/v_{0}$ in Eq. (7) converges, Weyl’s law implies the following:
   1. (iv)
-    The $\lambda_{n}$ can at most grow as $n^{2}$ for large $n$, i.e., $\lambda_{n}=O(n^{2})$.
-The last follows by Weyl’s law; see the SM [60] and, e.g., [78, 79].
-Simply put, it gives a semi-classical estimate of the number of eigenvalues below a given large $\lambda>0$ for Laplace-type operators.
+    $\lambda_{n}$ grows at most as $n^{2}$ for $n\gg 1$, i.e., $\lambda_{n}=O(n^{2})$;
+cf. the SM [74] and [97, 98].
 Unless otherwise stated, our SL problem will be regular.
@@
 The stated properties allow us to expand
-$\varphi(x)=\sum_{n=0}^{\infty}\varphi_{n}u_{n}(x),\qquad\Pi(x)=\sum_{n=0}^{\infty}\Pi_{n}w(x)u_{n}(x),$ (14)
-whereby we obtain $\bigl{[}\varphi_{n^{\vphantom{\prime}}},\Pi_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$ and can diagonalize $H_{\textrm{iTLL}}$ in Eq. (8). Indeed, the latter becomes
+$\varphi(x)=\sum_{n=0}^{\infty}\varphi_{n}u_{n}(x),\qquad\Pi(x)=\sum_{n=0}^{\infty}\Pi_{n}w(x)u_{n}(x)$ (14)
+with $\bigl[\varphi_{n^{\vphantom{\prime}}},\Pi_{n^{\prime}}\bigr]=\mathrm{i}\delta_{n,n^{\prime}}$ and thereby diagonalize $H_{\textrm{iTLL}}$ in Eq. (8). Indeed, the latter becomes
 $H_{\textrm{iTLL}}=\frac{1}{2\pi}\bigl[\pi\Pi_{0}\bigr]^{2}+\sum_{n=1}^{\infty}E_{n}a^{\dagger}_{n}a^{\vphantom{\dagger}}_{n}$ (15)
@@
 $E_{n}=\sqrt{\lambda_{n}},$ (16)
-where we expressed $\varphi_{n}=\sqrt{\pi/2E_{n}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right)$ and $\Pi_{n}=-\mathrm{i}\sqrt{E_{n}/2\pi}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ for $n\geq 1$ in bosonic operators $a^{\dagger}_{n}$, $a^{\vphantom{\dagger}}_{n}$ with $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$.
-Note that $\varphi_{0}$ (and zero modes in general) should be treated separately, but this will have no consequence here.
-In the SM [60], we also discuss SL operators (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq.
+where we expressed $\varphi_{n}=\sqrt{\pi/2E_{n}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right)$ and $\Pi_{n}=-\mathrm{i}\sqrt{E_{n}/2\pi}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ for $n\geq 1$ in bosonic operators $a^{\dagger}_{n}$, $a^{\vphantom{\dagger}}_{n}$ satisfying $\bigl[a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr]=\delta_{n,n^{\prime}}$ [99].
+[…]
+The SM [74] also examines operators of the form (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq.
 (8).
-The time evolution of any observable, such as $\varphi(x)$, will contain terms like $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$.
-Thus, as shown in the SM [60], to exhibit PWT as defined by Eq.
-(1), all $u_{n}(x)$ must have definite parity, i.e., be invariant under reflections $x\mapsto-x$ up to a sign, which must equal the value of each phase factor $\mathrm{e}^{\pm\mathrm{i}E_{n}T}$ for a common $T>0$.
-Moreover, the $u_{n}(x)$ have definite parity if and only if $v(x)$, $K(x)$ [and $q(x)$ if non-zero] are even, i.e., $H_{\textrm{iTLL}}$ is reflection symmetric.
-In conclusion, PWT boils down to requirements on the spectrum:
-An inhomogeneous TLL exhibits PWT if and only if $v(x)$ and $K(x)$ are even and all eigenvalues are of the form $E_{n}=\pi m_{n}/T$ for certain integers $m_{n}$.
-From Properties (i) and (ii) and since $\lambda_{0}=0$ here, these integers $m_{n}$ satisfy $0=m_{0}<m_{1}<\ldots$ with $m_{n}$ even (odd) if $n$ is even (odd).
-Moreover, from Property (iv), they scale as $m_{n}\sim n$ for large $n$, i.e., $m_{n}=n+o(n)$.
-Lastly, to probe PWT based on Eq.
-(5), we note that the results above allow one to compute correlation functions for inhomogeneous TLLs; see the SM [60].
-In particular, the ground-state $\varphi\varphi$ correlation function is
+The time evolution of any observable, such as $\varphi(x)$, contains terms like $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$.
+Thus, for Eq.
+(1) to hold for all states, each $u_{n}(x)$ must have definite parity, i.e., be invariant under reflections $x\mapsto-x$ up to a sign, which must equal each phase factor $\mathrm{e}^{\pm\mathrm{i}E_{n}T}$ for a common shortest possible $T>0$.
+Moreover, the $u_{n}(x)$ have definite parity if and only if $v(x)$, $K(x)$ [and $q(x)$ if nonzero] are even, i.e., $H_{\textrm{iTLL}}$ is reflection symmetric.
+This leads to a statement in terms of spectral requirements [74]: Inhomogeneous TLLs exhibit PWT as defined in Eq.
+(1) if and only if $v(x)$ and $K(x)$ are even and the eigenenergies are $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ so that [100]
+  1. (a)
+    $0=m_{0}<m_{1}<\ldots$ with each $m_{n}$ even or odd if $n$ is even or odd,
+  2. (b)
+    $m_{n}\sim n$ for large $n$, i.e., $m_{n}=n+o(n)$.
+To probe PWT based on Eq. (5), note that the results above allow one to compute correlations for inhomogeneous TLLs [74]. In particular, the ground-state $\varphi\varphi$ correlation function is [99]
 $\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle=\sum_{n=1}^{\infty}\frac{\pi}{2E_{n}}u_{n}(x)u_{n}(x^{\prime})\mathrm{e}^{-\mathrm{i}E_{n}(t-t^{\prime})}.$ (17)
@@
 (17) for $K(x)=1$, which, by bosonization, is related to $\log F(t)$ in Eq.
-(5), in that it encodes the same information; see the SM [60].
-Inverse Sturm-Liouville problem.—Reversing the logic, the PWT requirements on the spectrum yield an inverse spectral problem for SL operators with Neumann BCs [and $q(x)=0$]:
-Given a spectrum $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ as above, determine the even functions $v(x)$ and $K(x)$.
-A related prominent problem is the inverse scattering occurring in the theory of integrable PDEs, whose solution is given by the Gelfand-Levitan-Marchenko equation.
-However, that case is formulated on an infinite line, while here we are interested in a finite interval.
-Before a general discussion, we consider two examples:
-The inverse SL problem has a known solution when the spectrum is exactly linear: For $K(x)=K>0$ constant but $v(x)>0$ general, the SL operator is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq.
-(7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl{(}n\pi[L+2y]/2L\bigr{)}$ and $u_{0}(x)=\sqrt{v_{0}/L}$.
-This is precisely the conformal case of non-interacting right/left movers, enjoying PWT for any even $v(x)>0$.
-In the previous example, the SL operator was regular.
-Relaxing this condition by taking $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl{[}1-(2x/L)^{2}\bigr{]}^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$, we get an irregular problem for which Properties (i)(iv) remain valid; the BCs being that a solution is sufficiently well-behaved for Eq.
-(9) to hold.
-The eigenfunctions are given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$ [80, §18.3] up to normalizations, which have definite parity and satisfy Eq.
-(13) with $\lambda_{n}=(\pi v_{0}/L)^{2}n(n+2\alpha)$ for $v_{0}=2v/\pi$.
-Clearly, only the case $\alpha=0$ (Chebyshev polynomials of the first kind) exhibits PWT, while all $\alpha\neq 0$ do not, including Legendre polynomials ($\alpha=1/2$) and Chebyshev polynomials of the second kind ($\alpha=1$) [80, §18.7].
+(5), in that it encodes the same information [74].
+Inverse Sturm-Liouville problem—Reversing the logic, the spectral requirements for PWT yield an inverse problem for SL operators with Neumann BCs [101]: Given a spectrum $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ as above, determine the even functions $v(x)$, $K(x)$ [and $q(x)$ if nonzero].
+Before a general discussion, we consider two examples.
+First, the solution is known if the spectrum is exactly linear: For $K(x)=K$ constant but $v(x)$ general, then $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq.
+(7) with $E_{n}=\pi nv_{0}/L$ and $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl(n\pi[L+2y]/2L\bigr)$, $u_{0}(x)=\sqrt{v_{0}/L}$.
+This is the conformal case of uncoupled right and left movers, enjoying PWT for any even $v(x)>0$.
+Second, for $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl[1-(2x/L)^{2}\bigr]^{\alpha}$ with $\alpha>-1/2$ and $v,K>0$, we get an irregular problem possessing Properties (i)(iv) [102].
+The eigenfunctions are Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$ (up to normalizations), which have definite parity and satisfy Eq.
+(13) with $\lambda_{n}=(\pi v_{0}/L)^{2}n(n+2\alpha)$, $v_{0}=2v/\pi$; cf.
+Sec.
+18.3 in [103].
+Clearly, only $\alpha=0$ (Chebyshev polynomials of the first kind) exhibits PWT, while all $\alpha\neq 0$ do not, including $\alpha=1/2$ (Legendre polynomials) and $\alpha=1$ (Chebyshev polynomials of the second kind); cf.
+Sec.
+18.7 in [103].
 The respective $\varphi\varphi$ correlations are plotted in Fig.
 2.
-To solve the inverse SL problem in general, we write the operator (11) in Liouville normal form, as is always possible for regular problems [80, §1.13(viii)]: Let $\mathfrak{u}(y)\equiv u(x)\sqrt{K(x)}$ for $y=y(x)$ given by Eq.
+To solve the inverse SL problem in general, we write the operator (11) in Liouville normal form, as is always possible for regular problems; cf.
+Sec.
+1.13(viii) in [103]: Let $\mathfrak{u}(y)\equiv u(x)\sqrt{K(x)}$ for $y=y(x)$ in Eq.
 (7) and define
-Figure 2: Plot of $C(x,t)\equiv\operatorname{Re}\langle\Omega|\varphi(x,t)\varphi(-3L/8,0)|\Omega\rangle$ in Eq.
-(17) for eigenfunctions in the form of (a) Chebyshev polynomials of the first kind, which exhibit PWT, contrasted to (b) Legendre polynomials and (c) Chebyshev polynomials of the second kind, which do not.
-The upper row contains intensity plots of the correlations as functions of $x/L\in[-1/2,1/2]$ and $t/T\in[0,1]$, while the lower row compares their profiles at $t=0$ and $T$.
-Note the asymmetry in (b) and (c).
-The velocity $v(x)$ (same in all cases) and the respective Luttinger parameter $K(x)$ are shown in (d).
 $\hat{q}(y)\equiv-\frac{q(x)}{w(x)}+\frac{1}{\sqrt{K(x)}}v_{0}^{2}\partial_{y}^{2}\sqrt{K(x)}.$ (18)
@@
 $v_{0}^{2}\partial_{y}^{2}\mathfrak{u}(y)+[\lambda-\hat{q}(y)]\mathfrak{u}(y)=0$ (19)
-with unchanged BCs  [^4] For even $v(x)>0$, they are $v_{0}\partial_{y}\mathfrak{u}(y)-v(x)\Lambda(x)\mathfrak{u}(y)=0$ at $y=\pm L/2$ with $\Lambda(x)\equiv\partial_{x}\ln\sqrt{K(x)}$ [53], which imply Neumann BCs if $K^{\prime}(x)\big{|}_{x=\pm L/2}=0$.
-and if reflection symmetric in $x$, then so in $y$.
-Clearly, $\mathfrak{u}_{n}(y)$ have the same eigenvalues as the respective $u_{n}(x)$ and similarly form a complete orthonormal basis but with identity weight.
-Results in the literature often require two sets of eigenvalues for different BCs to get a unique solution to the inverse problem.
-However, one set suffices for reflection-symmetric SL operators and certain BCs, including ours.
-Indeed, [82] and [83, Theorem II] say that an even $\hat{q}(y)$ is uniquely determined by the spectrum for Neumann BCs.
-Thus, viewing $\hat{q}(y)$ as given and since $q(x)=0$, Eq.
-(18) becomes a second-order differential equation for $\sqrt{K(x(y))}$.
-By general uniqueness results [assuming $\hat{q}(y)$ is continuous] and since $K(x)$ must be even, this determines $K(x)$ up to changing coordinates (7) and an overall rescaling (discarding solutions that are not positive).
-The results [82, 83] do not reconstruct $\hat{q}(y)$.
-However, since we know that $K(x)=K$ gives an SL problem with $m_{n}=n$ and $\hat{q}(y)=0$, a constant Luttinger parameter is the unique solution for the inverse problem given an exactly linear spectrum.
-Beyond this special case, nothing a priori prevents the spectrum from deviating from the leading $n$ behavior as long as the requirements on the $m_{n}$ remain satisfied.
-This would then give rise to a class of possible bosonic theories that exhibit PWT, where, by uniqueness, $\hat{q}(y)$ is non-zero, thus $K(x)$ non-constant.
-While not usually analytically tractable, there are constructive approaches to find $\hat{q}(y)$ [84, 85, 86].
-Discussion.—We introduced the notion of PWT and showed that it is exhibited by one-particle excitations in inhomogeneous CFT whenever the velocity profile $v(x)>0$ is even.
-Breaking conformal invariance, by considering inhomogeneous TLLs with a Luttinger-parameter profile $K(x)>0$, we showed that PWT for general excitations amounts to an inverse spectral problem, which we solved.
-This continuum result for bosons is remarkably similar to earlier results: For spin chains with nearest neighbor hopping, the corresponding class of operators can be represented as Jacobi matrices, and the resulting inverse eigenvalue problem was studied [8, 87].
-As in the continuum, two sets of eigenvalues are generally needed for uniqueness [88, 86], while one set suffices for matrices symmetric with respect to the anti-diagonal (corresponding to reflection symmetry) [89, Theorem 3].
+with unchanged BCs [104], and if reflection symmetric in $x$, then so in $y$.
+Clearly, $\mathfrak{u}_{n}(y)$ have the same eigenvalues as the respective $u_{n}(x)$, and they similarly form a complete orthonormal basis but with identity weight.
+Two sets of eigenvalues for different BCs are usually required to uniquely solve an inverse SL problem.
+However, one set suffices for reflection-symmetric operators with certain BCs.
+Indeed, an even $\hat{q}(y)$ is uniquely determined by the spectrum for Neumann BCs [105, 106].
+Thus, taking $\hat{q}(y)$ as given, Eq.
+(18) [with $q(x)=0$] becomes a second-order differential equation for $\sqrt{K(x(y))}$.
+By uniqueness, this determines an even $K(x)$ up to changing coordinates as in Eq.
+(7) and an overall rescaling [assuming that $\hat{q}(y)$ is continuous and the solution $\sqrt{K(x)}$ is positive].
+The results [105, 106] do not reconstruct $\hat{q}(y)$, although there are constructive approaches [107, 108, 109].
+However, as $K(x)=K$ gives an SL problem with $m_{n}=n$ and $\hat{q}(y)=0$, a constant Luttinger parameter is the unique solution for the inverse problem given an exactly linear spectrum.
+Beyond this case, one may imagine that the spectrum could deviate from the leading $n$ behavior as long as the requirements on the $m_{n}$ are satisfied.
+By an exact Bohr-Sommerfeld/WKB quantization condition [110, 111, 112, 113], we show in the SM [74] that this depends on the regularity: If the system is regular also after unfolding the interval [L/2,L/2][-L/2,L/2] to the circle of twice the length [114], then PWT requires $\hat{q}(y)=0$.
+Thus, the conformal case of an exactly linear spectrum is the only possibility for such regular inhomogeneous TLLs to exhibit PWT.
+Discussion—We investigated the collective transfer of quantum information in continuous communication channels, introducing the notion of perfect wave transfer (PWT).
+Assuming conformal invariance, we showed that PWT is exhibited for one-particle excitations whenever the system is reflection symmetric.
+More generally, PWT with respect to arbitrary states amounts to an inverse spectral problem, studied in detail for inhomogeneous TLL theory—a prototypical model of 1+1D bosons with coupled right and left movers.
+Its solution shows that conformal invariance is essential for PWT if sufficient regularity is required.
+This approach is similar to results for spin chains with nearest-neighbor hopping: Their Hamiltonians can be represented as Jacobi matrices, yielding an inverse eigenvalue problem [9, 115] whose solution is unique [116, 109] given one set of eigenvalues for matrices symmetric with respect to the antidiagonal (reflection symmetry) [117].
+Remarkably, however, the continuum problem is more constrained due to asymptotic requirements, which imply that the spectrum must be exactly linear for inhomogeneous bosons to exhibit PWT under strong enough regularity conditions.
+This is not the case for discrete systems due to their intrinsic ultraviolet cutoff.
+Such a cutoff is also relevant for applications of TLL theory [49], and if present, there are no asymptotic requirements and more possibilities for PWT even for regular inhomogeneous TLLs.
+Moreover, as shown in the SM [74], PWT is also possible in inhomogeneous TLLs with fine-tuned mass terms, governed by a similar inverse spectral problem but with $q(x)\neq 0$.
 By bosonization, our results for inhomogeneous TLLs cover fermionic theories with position-dependent interactions [61].
-Indeed, even when $K(x)=K$ is constant, as long as $K\neq 1$, the theory describes an interacting quantum many-body system, albeit at low-energies.
-Lastly, we showed that an inhomogeneous TLL with constant $K(x)=K$ enjoys PWT for general excitations whenever $v(x)>0$ is even.
-Such a theory is an example of an inhomogeneous CFT.
-It would be interesting to study PWT for any number of excitations in general CFTs, and also extend to systems on the half or full infinite line.
-Acknowledgments.
-Acknowledgements.
-We are thankful to Pawel Caputa and Shinsei Ryu for motivating discussions on related topics.
+The essential property for these systems to exhibit PWT as in Eq.
+(1) was shown to be conformal invariance, assuming that the inhomogeneities are sufficiently regular, while it remains unclear what is possible if the regularity is relaxed.
+Consequently, PWT for all (not only one-particle) states holds for particular CFTs, such as inhomogeneous CFTs of free fermions and compactified bosons.
+It would be interesting to prove if PWT holds for any number of excitations in general CFTs and to extend to systems on the half or full infinite line.
+Our Letter can be seen as a first investigation of PWT for continuous quantum systems, focusing on universal low-energy descriptions within CFT and TLL theory.
+A second step would be to study PWT for specific gapless 1+1D models, including Lieb-Liniger [85].
+For those, our results apply at low energies, while full treatments require tailored solution methods, such as Bethe ansatz, but typically extended to inhomogeneous systems for PWT to be possible—a very difficult problem in itself.
+Generic interaction or curvature effects may also spoil PWT beyond the low-energy regime.
+Refinements may then be needed, either relaxing PWT to specific classes of observables or states, or reformulating it as almost PWT with fidelity $1-\varepsilon$ ($\varepsilon>0$) for the transfer of information.
+A concrete proposal for a general approach that builds on our present results would be to use nonlinear TLL theory [118, 119] to study such questions.
+Acknowledgments—We are thankful to Pawel Caputa, Shinsei Ryu, and Imke Schneider for motivating discussions.
 P.M.
-gratefully acknowledges financial support from the Wenner-Gren Foundations (Grant No.
+acknowledges financial support from the Wenner-Gren Foundations (Grant No.
 FT2022-0002).
 M.C.
-acknowledges financial support from the European Research Council (ERC Grant Agreement No.
+acknowledges financial support from the European Research Council (Grant No.
 818761), Villum Fonden via the QMATH Centre of Excellence (Grant No.
-10059) and the Novo Nordisk Foundation (Grant No.
-NNF20OC0059939 ‘Quantum for Life’), and thanks the NCCR SwissMAP of the Swiss National Science Foundation and the Section of Mathematics at the University of Geneva for their hospitality.
-Supplemental Material
+10059), and the Novo Nordisk Foundation (Grant No.
+NNF20OC0059939 “Quantum for Life”) and thanks the NCCR SwissMAP of the Swiss National Science Foundation and the Section of Mathematics at the University of Geneva for their hospitality.
+We also thank the organizers of the workshop “Integrability in Condensed Matter Physics and Quantum Field Theory” in 2023 at the SwissMAP Research Station in Les Diablerets, where this work was initiated.
+Supplemental Material for:
+Perfect Wave Transfer in Continuous Quantum Systems
 This supplemental material has three parts.
-Part A gives details for two-point correlations in inhomogeneous 1+1D CFT with open (Neumann) boundary conditions (BCs).
-Part B describes how to diagonalize general inhomogeneous Tomonaga-Luttinger liquids (TLLs) using Sturm-Liouville (SL) theory, leading to a formulation of perfect wave transfer (PWT) as an inverse SL problem, including extension to massive TLLs.
-Part C contains details for correlation functions in inhomogeneous TLLs.
+Part A gives details for two-point correlations in inhomogeneous 1+1D conformal field theory (CFT) with open boundary conditions (BCs).
+Part B describes how to diagonalize general inhomogeneous Tomonaga-Luttinger liquids (TLLs) using Sturm-Liouville (SL) theory, leading to the formulation of perfect wave transfer (PWT) as an inverse SL problem, including extension to “massive” TLLs and consequences inferred from Weyl’s law and the Bohr-Sommerfeld/WKB quantization condition.
+Part C contains details on correlations in inhomogeneous TLLs.
 Part A: Correlations in inhomogeneous 1+1D CFT with open boundaries
@@
 $H_{\textrm{iCFT}}\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,v(x)\bigl[T_{+}(x)+T_{-}(x)\bigr],$ (20)
-where $T_{\pm}(x)$ are the right/left-moving components of the stress-energy tensor in light-cone coordinates, whose Fourier transforms yield two commuting copies of the Virasoro algebra.
+where $T_{\pm}(x)$ are the right- (++) and left- (-) moving components of the stress-energy tensor in light-cone coordinates, whose Fourier transforms yield two commuting copies of the Virasoro algebra; see, e.g., [36].
 In position space, they satisfy the equal-time commutation relations
 $\bigl[T_{\pm}(x),T_{\pm}(x^{\prime})\bigr]=\mp 2\mathrm{i}\partial_{x}\delta(x-x^{\prime})T_{\pm}(x^{\prime})\pm\mathrm{i}\delta(x-x^{\prime})\partial_{x^{\prime}}T_{\pm}(x^{\prime})\pm\frac{c}{24\pi}\mathrm{i}\partial_{x}^{3}\delta(x-x^{\prime})$ (21)
-and $\bigl{[}T_{\pm}(x),T_{\mp}(x^{\prime})\bigr{]}=0$.
-Note that, at non-equal time, $T_{\pm}$ should be understood to only depend on the “right/left-moving coordinate” $x^{\mp}$; for now, we only view these as labels, distinguishing dependencies on right/left components.
+and $\bigl[T_{\pm}(x),T_{\mp}(x^{\prime})\bigr]=0$, where $c$ is the so-called central charge.
+Note that $T_{\pm}$ should be understood to only depend on the “right/left-moving coordinate” $x^{\mp}$, i.e., $T_{\pm}(x^{\mp})$; for now, we only view these as labels, distinguishing dependencies for right- or left-moving components.
 Besides the $T_{\pm}(x)$ fields, an important class of operators are Virasoro primary fields $\Phi(x^{-},x^{+})$, defined by how they behave under conformal transformations.
-The latter are given by orientation-preserving circle diffeomorphisms $f_{+}$, $f_{-}$ and primary fields transform as
+The latter are given by orientation-preserving circle diffeomorphisms $f_{+}$ and $f_{-}$, under which primary fields transform as
 $\Phi(x^{-},x^{+})\to f^{\prime}_{+}(x^{-})^{\Delta_{\Phi}^{+}}f^{\prime}_{-}(x^{+})^{\Delta_{\Phi}^{-}}\Phi(f_{+}(x^{-}),f_{-}(x^{+})),$ (22)
 where ($\Delta_{\Phi}^{+},\Delta_{\Phi}^{-}$) denote the field’s conformal weights.
-To impose open BCs at $x=\pm L/2$, as in the main text, we will assume that primary fields come in pairs $\Phi_{+}$, $\Phi_{-}$ with conformal weights satisfying $\Delta^{+}_{\Phi_{+}}=\Delta^{-}_{\Phi_{-}}$, $\Delta^{-}_{\Phi_{+}}=\Delta^{+}_{\Phi_{-}}$, since this allows one to couple excitations so that right movers are reflected to left movers when reaching $x=L/2$, and vice-versa at $x=-L/2$.
-To ease notation, we will use $\Delta^{+}_{\Phi}\equiv\Delta^{+}_{\Phi_{+}}$, $\Delta^{-}_{\Phi}\equiv\Delta^{-}_{\Phi_{+}}$ for each such pair.
+To impose open BCs at $x=\pm L/2$, as in the main text, we will assume that primary fields come in pairs $\Phi_{+}$, $\Phi_{-}$ with conformal weights satisfying $\Delta^{+}_{\Phi_{+}}=\Delta^{-}_{\Phi_{-}}$, $\Delta^{-}_{\Phi_{+}}=\Delta^{+}_{\Phi_{-}}$, since this allows one to couple excitations so that right movers are reflected to left movers when reaching $x=L/2$, and vice versa at x=L/2x=-L/2.
+For ease of notation, we will use $\Delta^{+}_{\Phi}\equiv\Delta^{+}_{\Phi_{+}}$, $\Delta^{-}_{\Phi}\equiv\Delta^{-}_{\Phi_{+}}$ for each such pair.
 Before continuing, we mention two examples:
   1. Example 1:
-    For the inhomogeneous free fermions in the main text, the components of the stress-energy tensor are $T_{\pm}(x)=(1/2)\bigl{[}\!\left.:\!\psi^{\dagger}_{\pm}(x)(\mp\mathrm{i}\partial_{x})\psi^{\vphantom{\dagger}}_{\pm}(x)\!:\right.+\mathrm{h.c.}\bigr{]}-\pi\big{/}12L^{2}$, and the primary fields of the theory are the fermionic fields $\psi_{\pm}(x)$ themselves, whose conformal weights are $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=1/2$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=0$.
+    For the inhomogeneous free fermions in the main text, the components of the stress-energy tensor are $T_{\pm}(x)=(1/2)\bigl[\!\left.:\!\hskip-0.5pt\psi^{\dagger}_{\pm}(x)(\mp\mathrm{i}\partial_{x})\psi^{\vphantom{\dagger}}_{\pm}(x)\hskip-0.5pt\!:\right.+\mathrm{h.c.}\bigr]-\pi\big/12L^{2}$, the central charge is $c=1$, and the primary fields of the theory are the fermionic fields $\psi_{\pm}(x)$ themselves, whose conformal weights are $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=1/2$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=0$.
   2. Example 2:
-    For an inhomogeneous TLL with constant Luttinger parameter $K(x)=K$, one has $T_{\pm}(x)=({1}/{4\pi K})\!\left.:\!\bigl{[}\pi\Pi(x)\mp K\partial_{x}\varphi(x)\bigr{]}^{2}\!:\right.-\pi\big{/}12L^{2}$, and the primary fields include vertex operators constructed as exponentials of the bosonic fields.
+    For an inhomogeneous TLL with constant Luttinger parameter $K(x)=K$, one has $T_{\pm}(x)=({1}/{4\pi K})\!\left.:\!\hskip-0.5pt\bigl[\pi\Pi(x)\mp K\partial_{x}\varphi(x)\bigr]^{2}\hskip-0.5pt\!:\right.-\pi\big/12L^{2}$, the central charge is $c=1$, and the primary fields include vertex operators constructed as exponentials of the bosonic fields.
 Applied to the Luttinger model [54, 55, 56], these vertex operators are (renormalized) fermionic fields $\psi_{\pm}$ with conformal weights $\Delta^{+}_{\psi_{+}}=\Delta^{-}_{\psi_{-}}=(1+K)^{2}/8K$, $\Delta^{-}_{\psi_{+}}=\Delta^{+}_{\psi_{-}}=(1-K)^{2}/8K$.
-(Clearly, free fermions are recovered for $K=1$.)
+(Free fermions are recovered by setting $K=1$.)
 As in the main text, open BCs can be incorporated by standard unfolding methods; cf. [75, 76, 65, 66, 48, 77]: Couple right- and left-moving excitations by identifying
@@
 for all pairs $\Phi_{+},\Phi_{-}$ of Virasoro primary fields.
-It follows that we can equivalently consider a chiral inhomogeneous CFT of only right movers on the interval $[-L/2,3L/2]$ of length $2L$ with velocity
-$\underline{v}(x)\equiv\begin{cases}v(x)&\text{for }x\in[-L/2,L/2],\\ v(L-x)&\text{for }x\in[L/2,3L/2]\end{cases}$ (24)
+It follows that we can equivalently consider a chiral inhomogeneous CFT of only right movers on the interval [L/2,3L/2][-L/2,3L/2] of length $2L$ with velocity
+$\underline{v}(x)\equiv\begin{cases}v(x)&\text{for }x\in[-L/2,L/2],\\ v(L-x)&\text{for }x\in[L/2,3L/2],\end{cases}$ (24)
 and with every quantity satisfying periodic BCs. Note that the Hamiltonian (20) can then be written as
@@
 $|\Phi_{j}\rangle\equiv\int_{-L/2}^{L/2}\mathrm{d}x\,\left(\mathrm{e}^{-\mathrm{i}kx}\xi^{+}_{j}(x)\Phi_{+}^{\dagger}(x,x)+\mathrm{e}^{\mathrm{i}kx}\xi^{-}_{j}(x)\Phi_{-}^{\dagger}(x,x)\right)|\Omega\rangle\qquad(j=1,2)$ (30)
-for smooth enough functions $\xi^{\pm}_{j}(x)$, interpreted as distributions or “wave packets” of right/left-moving excitations in each state, and some $k\in\mathbb{R}$ in order to accommodate for $k_{\textrm{F}}$ in Eq.
+for smooth enough functions $\xi^{\pm}_{j}(x)$, interpreted as “waves” of right- or left-moving excitations in each state, and some $k\in\mathbb{R}$ in order to accommodate for $k_{\textrm{F}}$ in Eq.
 (3).
-(The latter will have no effect on PWT, and one might as well set $k=0$.) As in the main text, this is a restricted class of states which only allow for one-particle excitations, while we eased other restrictions in the sense that the excitations are not necessarily fermionic and the distributions of right/left movers can be different.
+(The latter will have no effect on PWT, and one might as well set $k=0$.) As in the main text, this is a restricted class of states which only allow for one-particle excitations, while we eased other restrictions in the sense that the excitations are not necessarily fermionic and the distributions of right and left movers in the “waves” can be different.
 To probe PWT for one-particle excitations, we seek to compute
@@
 $F(t)\equiv\langle\Phi_{2}|\mathrm{e}^{-\mathrm{i}H_{\textrm{iCFT}}t}|\Phi_{1}\rangle=\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\langle\Omega|\left(\mathrm{e}^{\mathrm{i}kx_{2}}\overline{\xi^{+}_{2}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(x_{2},x_{2})+\mathrm{e}^{-\mathrm{i}kx_{2}}\overline{\xi^{-}_{2}(x_{2})}\Phi_{-}^{\vphantom{\dagger}}(x_{2},x_{2})\right)\mathrm{e}^{-\mathrm{i}H_{\textrm{iCFT}}t}\\ \times\left(\mathrm{e}^{-\mathrm{i}kx_{1}}\xi^{+}_{1}(x_{1})\Phi_{+}^{\dagger}(x_{1},x_{1})+\mathrm{e}^{\mathrm{i}kx_{1}}\xi^{-}_{1}(x_{1})\Phi_{-}^{\dagger}(x_{1},x_{1})\right)|\Omega\rangle,$ (32)
-in order to eventually compare $|F(t)|$ with $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for specular waves, $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$. Using the unfolding relations (23), we obtain
+in order to eventually compare $|F(t)|$ with $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for specular waves, $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$. Using the unfolding relations in Eq. (23), we obtain
 $\langle\Phi_{j}|\Phi_{j}\rangle=\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\langle\Omega|\left(\mathrm{e}^{\mathrm{i}kx_{2}}\overline{\xi^{+}_{j}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(x_{2},x_{2})+\mathrm{e}^{-\mathrm{i}kx_{2}}\overline{\xi^{-}_{j}(x_{2})}\Phi_{+}^{\vphantom{\dagger}}(L-x_{2},L-x_{2})\right)\\ \times\left(\mathrm{e}^{-\mathrm{i}kx_{1}}\xi^{+}_{j}(x_{1})\Phi_{+}^{\dagger}(x_{1},x_{1})+\mathrm{e}^{\mathrm{i}kx_{1}}\xi^{-}_{j}(x_{1})\Phi_{+}^{\dagger}(L-x_{1},L-x_{1})\right)|\Omega\rangle$ (33)
@@
 where we used that $H_{\textrm{iCFT}}|\Omega\rangle=C|\Omega\rangle$ for some real constant $C$ that encodes the Casimir energy.
-(This only gives rise to an overall phase that one could safely ignore.) Clearly, as stated in the main text, the above amounts to evaluating two-point correlations for primary fields, specifically those of right movers in a chiral CFT on $[-L/2,3L/2]$.
-First, as a preliminary step, for primary fields with periodic boundary conditions on the system with length $2L$, conformal invariance implies that the undeformed two-point correlations are
+(This only gives rise to an overall phase that one could safely ignore.) Clearly, as stated in the main text, the above amounts to evaluating and comparing two-point correlations for primary fields, specifically those of right movers in a chiral CFT on [L/2,3L/2][-L/2,3L/2].
+First, as a preliminary step, for primary fields in a system of length $2L$ with periodic BCs, conformal invariance implies that the (undeformed) two-point correlations are of the form
 $\langle\Omega_{0}|\Phi^{\vphantom{\dagger}}_{+}(x_{2}^{-},x_{2}^{+})\Phi^{\dagger}_{+}(x_{1}^{-},x_{1}^{+})|\Omega_{0}\rangle=\frac{1}{2\pi}G^{+}_{\Phi\Phi}(x_{2}^{-}-x_{1}^{-})^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}(x_{2}^{+}-x_{1}^{+})^{2\Delta^{-}_{\Phi}},$ (35)
-where $G^{\pm}_{\Phi\Phi}(x)$ are $2L$-periodic functions.
-For example, as can be seen from, e.g., [55], the correlations for right-moving free fermions (cf.
-Example 1 above) on the system with length $2L$ and periodic BCs are
+where $G^{\pm}_{\Phi\Phi}(x)$ are $2L$-periodic functions. For example, as can be seen from, e.g., [63, 67], the two-point correlations for right-moving free fermions (cf. Example 1 above) are
 $\langle\Omega_{0}|\psi^{\vphantom{\dagger}}_{+}(x_{2})\psi^{\dagger}_{+}(x_{1})|\Omega_{0}\rangle=\frac{1}{2\pi}G^{+}_{\textrm{FF}}(x_{2}-x_{1})$ (36)
@@
 $G^{+}_{\textrm{FF}}(x)\equiv\frac{\mathrm{i}\pi\exp\bigl(-\mathrm{i}\frac{\pi}{2L}x\bigr)}{2L\sin\bigl(\frac{\pi}{2L}[x+\mathrm{i}0^{+}]\bigr)},$ (37)
-where we used that $\Delta^{+}_{\psi}=1/2$, $\Delta^{-}_{\psi}=0$. Similarly, for the Luttinger model (cf. Example 2 above), the correlations for the (renormalized) right-moving fermionic fields are
+where we used that $\Delta^{+}_{\psi}=1/2$, $\Delta^{-}_{\psi}=0$.
+Similarly, for the Luttinger model (cf.
+Example 2 above), the two-point correlations for the (renormalized) right-moving fermionic fields are
 $\langle\Omega_{0}|\psi^{\vphantom{\dagger}}_{+}(x_{2}^{-},x_{2}^{+})\psi^{\dagger}_{+}(x_{1}^{-},x_{1}^{+})|\Omega_{0}\rangle=\frac{1}{2\pi}G^{+}_{\textrm{FF}}(x_{2}^{-}-x_{1}^{-})^{(1+K)^{2}/4K}G^{+}_{\textrm{FF}}(x_{1}^{+}-x_{2}^{+})^{(1-K)^{2}/4K},$ (38)
@@
 In addition, the action of $U_{+}(\underline{f})$ from the expression (29) for $|\Omega\rangle$ also leads to a conformal transformation of the fields.
-It follows from this and the properties (27) that
+It follows from this and the properties in Eq.
+(27) that
 $\langle\Phi_{j}|\Phi_{j}\rangle=\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl(&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{j}(x_{2})}\xi^{+}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl(\underline{f}(x_{2})-\underline{f}(x_{1})\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(\underline{f}(x_{2})-\underline{f}(x_{1})\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{j}(x_{2})}\xi^{-}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl(\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{j}(x_{2})}\xi^{+}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{j}(x_{2})}\xi^{-}_{j}(x_{1})G^{+}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})+\underline{f}(x_{1})\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})+\underline{f}(x_{1})\bigr)^{2\Delta^{-}_{\Phi}}\biggr)\end{aligned}$ (40)
 and
-$F(t)=\mathrm{e}^{-\mathrm{i}Ct}\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl{(}&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{v}_{0}t\bigr{)}^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl{(}-\underline{f}(x_{2})+\underline{f}(x_{1})+\underline{v}_{0}t\bigr{)}^{2\Delta^{-}_{\Phi}}\biggr{)},\end{aligned}$ (41)
+$\hskip-7.11317ptF(t)=\mathrm{e}^{-\mathrm{i}Ct}\frac{1}{2\pi}\int_{-L/2}^{L/2}\mathrm{d}x_{1}\,\int_{-L/2}^{L/2}\mathrm{d}x_{2}\,\underline{f}^{\prime}(x_{2})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{+}_{\Phi}}\underline{f}^{\prime}(x_{2})^{\Delta^{-}_{\Phi}}\underline{f}^{\prime}(x_{1})^{\Delta^{-}_{\Phi}}\\ \begin{aligned} \times\biggl(&\mathrm{e}^{\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(\underline{f}(x_{2})-\underline{f}(x_{1})-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{+}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{f}(L)+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}+x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{+}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})-\underline{f}(x_{1})+\underline{f}(L)+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\\ &+\mathrm{e}^{-\mathrm{i}k(x_{2}-x_{1})}\overline{\xi^{-}_{2}(x_{2})}\xi^{-}_{1}(x_{1})G^{+}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})+\underline{f}(x_{1})-\underline{v}_{0}t\bigr)^{2\Delta^{+}_{\Phi}}G^{-}_{\Phi\Phi}\bigl(-\underline{f}(x_{2})+\underline{f}(x_{1})+\underline{v}_{0}t\bigr)^{2\Delta^{-}_{\Phi}}\biggr),\end{aligned}$ (41)
 where we used that $\underline{f}\bigl(\underline{x}_{t}^{\pm}(x)\bigr)=\underline{f}(x)\pm\underline{v}_{0}t$ and $\underline{f}\bigl(\underline{x}_{t}^{\pm}(L-x)\bigr)=\underline{f}(L)-\underline{f}\bigl(\underline{x}_{t}^{\mp}(x)\bigr)=\underline{f}(L)-\underline{f}(x)\pm\underline{v}_{0}t$.
@@
 (42) and (40), it follows that $|F(t)|$ for certain $t$ equals $|\langle\Phi_{1}|\Phi_{1}\rangle|$ for all $\xi^{\pm}_{1}(x)$ if and only if $\underline{v}(x)$ is even, meaning $\underline{f}(x)$ must be odd.
-This in turn implies $\underline{f}(L)=L$ and $\underline{v}_{0}=v_{0}$ given by Eq.
+This in turn implies that $\underline{f}(L)=L$ and $\underline{v}_{0}=v_{0}$ given by Eq.
 (7), which means that equality holds for all $t=(2m+1)L/v_{0}$ due the $2L$ periodicity of $G^{\pm}_{\Phi\Phi}(x)$.
 [Moreover, for $\xi^{\pm}_{2}(x)=\xi^{\mp}_{1}(-x)$, one has $|\langle\Phi_{1}|\Phi_{1}\rangle|=|\langle\Phi_{2}|\Phi_{2}\rangle|$ when $v(x)$ is even.] Since $\underline{v}(x)=v(x)$ for all $x\in[-L/2,L/2]$, this proves the claim that one-particle PWT is exhibited for all even inhomogeneous CFTs.
-As a last remark, one can unfold in a different way to take the spin/statistics of the primary fields into account, so that, e.g., one gets anti-periodic BCs for fermions on the system of size $2L$, and twisted BCs for general anyonic fields.
-This can be used to obtain simple explicit formulas for the two-point correlations of any primary fields.
-However, since we are interested in recombinations of waves traveling in different directions, it is necessary to identify so as to always have periodic BCs.
-In any case, for our purposes, all we needed was the spacetime dependence of the correlations, which are fixed on general grounds in CFT, circumventing any need for explicit formulas.
+As a last remark, it is possible to unfold in different ways to incorporate the statistics of the primary fields.
+E.g., so that, for the system of length $2L$, one gets antiperiodic BCs for fermions and twisted BCs for general anyonic fields.
+However, since we are interested in recombinations of waves traveling in different directions, we found it necessary to identify right and left movers so as to always have periodic BCs.
 Part B: Spectral properties of inhomogeneous TLLs
-In this part, we first show that reflection symmetry is a necessary condition for PWT for any theory; here under the assumption that the Hamiltonian has pure point spectrum.
-Second, we give the details for how to use SL theory to diagonalize the inhomogeneous TLL Hamiltonian (8) and to formulate the PWT requirements on the spectrum.
-Third, we review some details on Weyl’s law applied to SL operators and, fourth, extend the previous discussion to TLL Hamiltonians with a mass term.
-Fifth, we comment on the Liouville normal form used to state the results for the inverse SL problem in the main text.
+In this part, we first show that reflection symmetry is a necessary condition for PWT in any system; here, under the assumption that the Hamiltonian has pure point spectrum.
+Second, we give the details for how to use SL theory to diagonalize the inhomogeneous TLL Hamiltonian (8) and how to formulate the PWT requirements on the spectrum.
+Third and fourth, respectively, we review certain details for Weyl’s law and the Bohr-Sommerfeld/WKB quantization condition applied to SL operators, and, fifth, extend the previous discussion to TLL Hamiltonians with a mass term.
+Sixth, we comment on the Liouville normal form used to state the results for the inverse SL problem in the main text.
 Necessity of reflection symmetry
@@
 (7) converges].
-Thus, we first consider a general Hamiltonian $H$ with pure point spectrum and show that for such a system to exhibit PST/PWT, it is necessary that $H$ commutes with the parity operator $\mathcal{P}$, which maps $x\mapsto-x$.
+Thus, we first consider a general Hamiltonian $H$ with pure point spectrum and show that for such a system to exhibit PWT or PST, it is necessary that $H$ commutes with the parity operator $\mathcal{P}$, which maps $x\mapsto-x$.
 In other words, we will show that reflection symmetry is a necessary condition.
@@
 (We recall that $\mathcal{P}$ is a unitary operator satisfying $\mathcal{P}^{2}=1$, meaning that $\mathcal{P}^{-1}=\mathcal{P}^{\dagger}=\mathcal{P}$.) As usual, define $F(t)\equiv\langle\Phi_{2}|\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle=\langle\Phi_{1}|\mathcal{P}\mathrm{e}^{-\mathrm{i}Ht}|\Phi_{1}\rangle$.
-For general PST/PWT, we require that there exists $T>0$ such that $|F(T)|=|\langle\Phi_{1}|\Phi_{1}\rangle|$ for any state $|\Phi_{1}\rangle$.
-In particular, this must hold for any energy eigenstate $|n\rangle$, implying $|\langle n|\mathcal{P}|n\rangle|=1$.
-If $|n\rangle|$ is not an eigenstate of $\mathcal{P}$, it is straightforward to show that this equality cannot hold.
-Consequently, all energy eigenstates must also be eigenstates of $\mathcal{P}$, which implies that $[\mathcal{P},H]=0$.
+Full PWT, as defined in Eq.
+(1), requires that there exists a common shortest possible time $T>0$ such that $|F(T)|=|\langle\Phi_{1}|\Phi_{1}\rangle|$ for all states $|\Phi_{1}\rangle$.
+In particular, this must hold for any energy eigenstate $|n\rangle$, implying that $|\langle n|\mathcal{P}|n\rangle|=1$.
+If $|n\rangle$ is not an eigenstate of $\mathcal{P}$, then it is straightforward to show that this equality cannot hold.
+Consequently, all energy eigenstates must also be eigenstates of $\mathcal{P}$, which implies that $\mathcal{P},H]=0$.
 Inhomogeneous TLLs and SL theory
-We now turn to inhomogeneous TLLs studied in the main text. Inserting the expansions (14) into Eq. (10) yields
-$H_{\textrm{iTLL}}=\frac{1}{2\pi}[\pi\Pi_{0}]^{2}+\frac{1}{2\pi}\sum_{n=1}^{\infty}\left.:\!\left([\pi\Pi_{n}]^{2}+\lambda_{n}\varphi_{n}^{2}\right)^{2}\!:\right.,$ (43)
-where $\varphi_{0}$ is absent (when $E_{0}=0$), no different from the standard case with constant $K(x)$ [and $v(x)$]. For $n\geq 1$, let
+We now turn to inhomogeneous TLL theory studied in the main text. Inserting the expansions in Eq. (14) into the Hamiltonian (10) and using Eq. (13) yield
+$H_{\textrm{iTLL}}=\frac{1}{2\pi}[\pi\Pi_{0}]^{2}+\frac{1}{2\pi}\sum_{n=1}^{\infty}\left.:\!\hskip-0.5pt\Bigl([\pi\Pi_{n}]^{2}+\lambda_{n}\varphi_{n}^{2}\Bigr)\hskip-0.5pt\!:\right.$. (43)
+Here, $\varphi_{0}$ is absent (when $E_{0}=0$), no different from the standard case with constant $K(x)$ [and $v(x)$]. For $n\geq 1$, let
 $\varphi_{n}=\sqrt{\frac{\pi}{2}}\frac{1}{\lambda_{n}^{1/4}}\left(a^{\vphantom{\dagger}}_{n}+a^{\dagger}_{n}\right),\qquad\Pi_{n}=-\mathrm{i}\frac{\lambda_{n}^{1/4}}{\sqrt{2\pi}}\left(a^{\vphantom{\dagger}}_{n}-a^{\dagger}_{n}\right)$ (44)
-in terms of bosonic creation and annihilation operators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$, satisfying $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$, as in the main text.
+in terms of bosonic creation and annihilation operators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$ satisfying $\bigl[a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr]=\delta_{n,n^{\prime}}$ and $a_{n}|\Omega\rangle=0$, where $|\Omega\rangle$ denotes the ground state of $H_{\textrm{iTLL}}$; cf.
+Sec.
+3.2 in [37].
 Inserting Eq.
@@
 $\varphi(x,t)=\sum_{n=1}^{\infty}\sqrt{\frac{\pi}{2E_{n}}}\left(a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}+a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}\right)u_{n}(x),\qquad\Pi(x,t)=\sum_{n=1}^{\infty}(-\mathrm{i})\sqrt{\frac{E_{n}}{2\pi}}\left(a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}-a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}\right)w(x)u_{n}(x),$ (45)
-where we omitted the zero modes (as they have to be treated separately when $E_{0}=0$) and used $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}$ and $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\dagger}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}$.
+where we omitted the zero modes (they must be treated separately when $E_{0}=0$, but in that case they will have no consequence for our discussion of PWT) and used $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}$ and $\mathrm{e}^{\mathrm{i}H_{\textrm{iTLL}}t}a^{\dagger}_{n}\mathrm{e}^{-\mathrm{i}H_{\textrm{iTLL}}t}=a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}$.
 Generally, any observable $\mathcal{O}(x,t)$ in this algebra will contain contributions of the form $a^{\dagger}_{n}\mathrm{e}^{\mathrm{i}E_{n}t}u_{n}(x)$ and $a^{\vphantom{\dagger}}_{n}\mathrm{e}^{-\mathrm{i}E_{n}t}u_{n}(x)$.
 Now, in order for PWT to be exhibited, Eq.
-(1) must hold for any state, which means that it holds true without expectation values, i.e., $\mathcal{O}(x,0)$ must equal $\mathcal{O}(-x,T)$ at some later time $T>0$ as an operator identity.
-Taking $\mathcal{O}=\varphi$ and using the linear independence of the generators $a^{\vphantom{\dagger}}_{n}$ and $a^{\dagger}_{n}$, Eq.
-(45) yields
+(1) must hold for any state in the Hilbert space, which means that it must hold without taking expectation values.
+I.e., $\mathcal{O}(x,0)$ must equal $\mathcal{O}(-x,T)$ at some later time $T>0$ as an operator identity.
+Taking $\mathcal{O}=\varphi$ and using the linear independence of the generators $a^{\dagger}_{n}$ and $a^{\vphantom{\dagger}}_{n}$, Eq.
+(45) implies that
 $u_{n}(x)=u_{n}(-x)\mathrm{e}^{\pm\mathrm{i}E_{n}T}\quad\forall n=0,1,\ldots$ (46)
-for a common $T>0$, where we included the trivial case $n=0$ [recall that $u_{0}(x)=\textrm{const}$ and $E_{0}=0$ when $q(x)=0$ in Eq.
+for a common shortest possible $T>0$, where we included the trivial case $n=0$ [recall that $u_{0}(x)=\textrm{const}$ and $E_{0}=0$ when $q(x)=0$ in Eq.
 (11)].
 Since $\mathcal{P}u_{n}(x)=u_{n}(-x)$, Eq.
-(46) implies that each $u_{n}(x)$ must be an eigenfunction of the parity operator $\mathcal{P}$, whose eigenvalues are $+1$ or $-1$ (since $\mathcal{P}^{2}=1$).
-It follows that all eigenfunctions must have definite parity, and the only possibility for PWT to hold is that the corresponding eigenenergies satisfy $E_{n}T=\pi m_{n}$ for integers $m_{n}$ for each $n$, with $m_{n}$ even (odd) if $u_{n}(x)$ has even (odd) parity.
-Finally, it is straightforward to show the following statement: The eigenfunctions $u_{n}(x)$ of a regular SL operator $\mathcal{A}$ have definite parity if and only if $[\mathcal{A},\mathcal{P}]=0$.
+(46) implies that each $u_{n}(x)$ must be an eigenfunction of the parity operator $\mathcal{P}$, whose eigenvalues are +1+1 or 1-1 (since $\mathcal{P}^{2}=1$).
+It follows that all eigenfunctions must have definite parity, and the only possibility for PWT to hold is that the corresponding eigenenergies satisfy $E_{n}T=\pi m_{n}$ for integers $m_{n}$, with $m_{n}$ even (odd) if $u_{n}(x)$ has even (odd) parity.
+Finally, it is straightforward to show the following statement: The eigenfunctions $u_{n}(x)$ of a regular SL operator $\mathcal{A}$ have definite parity if and only if $\mathcal{A},\mathcal{P}]=0$.
 For our operator in Eqs.
 (11)–(12), this means that both $v(x)$ and $K(x)$ must be even functions (using that they must be positive to exclude the possibility that both are instead odd).
-Given the above, together with Properties (i)(iv) for regular SL operators stated in the main text, we have derived our main result on the spectral requirements for an inhomogeneous TLL to exhibit PWT.
-Moreover, as stated in the main text, reversing the logic, namely, given a spectrum that satisfy the spectral properties for PWT to be exhibited, our formulation of PWT for inhomogeneous TLLs as an inverse SL problem follows.
+Given the above, together with Properties (i)(iv) for regular SL operators stated in the main text, we have derived our statement in terms of spectral requirements for an inhomogeneous TLL to exhibit PWT.
+Moreover, as stated in the main text, reversing the logic, i.e., assuming we are given a spectrum that satisfies the spectral requirements for PWT, we obtain our formulation of PWT for inhomogeneous TLLs as an inverse problem.
 Weyl’s law for SL operators
-Weyl’s law applied to SL operators, see, e.g., [78, 79], states that the number $N(\lambda)$ of eigenvalues below a given $\lambda>0$ asymptotically behaves as
+Weyl’s law applied to SL operators of the general form (11) states that the number $N(\lambda)$ of eigenvalues below a given $\lambda>0$ asymptotically behaves as
 $\frac{N(\lambda)}{\sqrt{\lambda}}\sim\frac{1}{\pi}\int_{-L/2}^{L/2}\mathrm{d}x\,\sqrt{\frac{w(x)}{p(x)}}$ (47)
-for large $\lambda$.
-This holds under certain technical conditions on the coefficients $w(x)$, $p(x)$ in Eq.
-(11) [which are satisfied assuming our $v(x)$, $K(x)$ are positive and smooth functions].
+for large $\lambda$; see, e.g., [97, 98].
+This holds under certain technical conditions on the functions $w(x)$, $p(x)$, and $q(x)$ in Eq.
+(11), which are satisfied assuming our $v(x)$ and $K(x)$ are positive and smooth functions [if $q(x)$ is nonzero, then it is also assumed to be smooth].
 Using Eqs.
@@
 (7).
-Thus, assuming that the integral defining $1/v_{0}$ converges, then this gives the stated asymptotic behavior, $E_{n}=\pi v_{0}n/L+o(n)$, in the main text.
-(Another possibility would be if the integral diverges, in which case $E_{n}\sim\textrm{const}\times n^{\gamma}$ for $\gamma<1$, but this is inconsistent with that the eigenenergies must be proportional to integers $m_{n}$ with a common proportionality constant in order for the theory to exhibit PWT.)
+Thus, assuming that the integral defining $1/v_{0}$ converges, this gives the stated asymptotic behavior, $E_{n}=\pi v_{0}n/L+o(n)$, in the main text.
+(Another possibility would be if the integral diverges, in which case $E_{n}\sim\textrm{const}\times n^{\gamma}$ for $\gamma<1$, but this is inconsistent with that the eigenenergies must be proportional to integers $m_{n}$ with a common proportionality constant, in order for the theory to exhibit PWT.)
+Exact Bohr-Sommerfeld/WKB quantization condition
+A stronger version of Weyl’s law for our class of operators can be deduced from an exact generalization of the usual formulation of the Bohr-Sommerfeld/WKB quantization condition.
+To state this condition, we consider our Sturm-Liouville problem written on Liouville normal form (19):
+$\bigl[-\partial_{y}^{2}+V(y)\bigr]\mathfrak{u}(y)=\Lambda\mathfrak{u}(y)$ (49)
+on the interval $I\equiv[-L/2,L/2]\ni y$, where we introduced $V(y)\equiv\hat{q}(y)/v_{0}^{2}$ and $\Lambda\equiv\lambda/v_{0}^{2}$ and recall that we impose Neumann BCs.
+The first terms of the exact quantization condition for the $n$th eigenvalue $\Lambda_{n}$ can then be stated as
+$\phi(\Lambda_{n})=\pi n$ (50)
+with
+$\phi(\Lambda)=\phi_{0}(\Lambda)+\phi_{1}(\Lambda)+O(\Lambda^{-5/2}),\vskip-7.11317pt$ (51)
+$\phi_{0}(\Lambda)\equiv\int_{I}\mathrm{d}y\,\sqrt{\Lambda-V(y)},\qquad\phi_{1}(\Lambda)\equiv-\frac{1}{4}\frac{\mathrm{d}}{\mathrm{d}\Lambda}\int_{I}\mathrm{d}y\,\frac{V^{\prime\prime}(y)}{\sqrt{\Lambda-V(y)}}.$ (52)
+This follows from [110, 111, 112] with the difference that we have no soft turning points but instead a finite interval with Neumann BCs: There is thus no Maslov index, and so the right-hand side of Eq.
+(50) is $\pi n$ instead of $\pi(n+1/2)$; cf.
+Sec.
+2.4 in [113].
+First, we consider $\phi_{1}(\Lambda)$. Using the expansion $1/\sqrt{\Lambda-V(y)}=\Lambda^{-1/2}\bigl[1+V(y)/2\Lambda+O(\Lambda^{-2})\bigr]$ for large $\Lambda$, we obtain
+$\int_{I}\mathrm{d}y\,\frac{V^{\prime\prime}(y)}{\sqrt{\Lambda-V(y)}}=\frac{1}{\sqrt{\Lambda}}\int_{I}\mathrm{d}y\,V^{\prime\prime}(y)+O(\Lambda^{-3/2}),$ (53)
+which implies that
+$\phi_{1}(\Lambda)=\frac{L}{8\Lambda^{3/2}}\Delta+O(\Lambda^{-5/2})$ (54)
+with
+$\Delta\equiv\frac{V^{\prime}(L/2)-V^{\prime}(-L/2)}{L}.$ (55)
+Recall that $V(y)$ must be an even function on $I$ (for the system to exhibit PWT) and thus $V^{\prime}(y)$ will be odd.
+Second, we consider $\phi_{0}(\Lambda)$.
+Using the expansion $\sqrt{\Lambda-V(y)}=\Lambda^{1/2}\bigl[1-V(y)/2\Lambda-V(y)^{2}/8\Lambda^{2}+O(\Lambda^{-3})\bigr]$ for large $\Lambda$, we find
+$\phi_{0}(\Lambda)=\sqrt{\Lambda}L\left(1-\frac{1}{2\Lambda}a_{1}-\frac{1}{8\Lambda^{2}}a_{2}\right)+O(\Lambda^{-5/2})$ (56)
+with
+$a_{1}\equiv\frac{1}{L}\int_{I}\mathrm{d}y\,V(y),\qquad a_{2}\equiv\frac{1}{L}\int_{I}\mathrm{d}y\,V(y)^{2},$ (57)
+which can be interpreted as the first two moments of the potential $V(y)$.
+Inserting the above into Eqs. (50)–(52) yields
+$\sqrt{\Lambda_{n}}\left(1-\frac{1}{2\Lambda_{n}}a_{1}-\frac{1}{8\Lambda_{n}^{2}}[a_{2}-\Delta]+O(\Lambda_{n}^{-3})\right)=\frac{\pi n}{L}.$ (58)
+Since Eq.
+(16) implies $\Lambda_{n}=\lambda_{n}/v_{0}^{2}=(E_{n}/v_{0})^{2}$ and since $E_{n}=\pi m_{n}/T$ for integers $m_{n}$ satisfying the conditions (a)(b) on page (a) of the main text, the quantization condition becomes
+$m_{n}\left(1-\frac{L^{2}}{2\pi^{2}m_{n}^{2}}a_{1}-\frac{L^{4}}{8\pi^{4}m_{n}^{4}}[a_{2}-\Delta]+O(m_{n}^{-6})\right)=n,$ (59)
+where we used that $v_{0}=L/T$ by definition of $T$ as the common shortest possible time (for all modes) for the system to exhibit PWT.
+Considering the difference $m_{n+1}-m_{n}$, it follows from Eq.
+(59) that $m_{n+1}-m_{n}=1$ for all $n>n_{0}$ for some sufficiently large positive integer $n_{0}$.
+The solution of this recursion formula is $m_{n}=n+a$ for some $a\in\mathbb{R}$ for $n>n_{0}$.
+In other words, $m_{n}=n\bigl[1+O(n^{-1})\bigr]$ for large $n$, which compared with Eq.
+(59) implies
+$a_{1}=0,\qquad a_{2}-\Delta=0.$ (60)
+I.e., the potential $V(y)$ must have a vanishing mean, which implies that $a_{2}$ is the variance, which must equal $\Delta$.
+[Note that Eqs.
+(55) and (57) then imply that $V^{\prime}(L/2)\geq 0$, since $\Delta=a_{2}\geq 0$ could otherwise not hold.]
+The interval $I$ with Neumann BCs can be unfolded to [L/2,3L/2][-L/2,3L/2] of length $2L$ with periodic BCs (similar to Part A), which in turn can be viewed as the circle $S^{1}_{2L}$.
+If we impose that $V(y)$ must be regular not only before but also after this unfolding procedure, then the only possibility is $V^{\prime}(L/2)=0$, since $V(y)$ must be even, from which it follows that $\Delta=0$ in Eq.
+(55).
+Consequently, both the mean $a_{1}$ and the variance $a_{2}$ of the potential $V(y)$ vanish in that case, which then implies that $\hat{q}(y)=v_{0}^{2}V(y)$ must be zero (almost everywhere).
+From Eq.
+(18) with $q(x)=0$ and reflection symmetry, it follows that $K(x)$ must be constant for PWT to be exhibited in an inhomogeneous TLL with the assumption of regularity after unfolding to $S^{1}_{2L}$.
+However, if such regularity is not assumed, then the quantization condition (59) does not prevent $\hat{q}(y)=v_{0}^{2}V(y)$ from being nonzero, which in turn corresponds to a nonconstant $K(x)$.
 “Massive” inhomogeneous TLLs
 For massless TLL bosons, as discussed above and in the main text, given a spectrum satisfying the PWT requirements, we saw that $K(x)$ can be found uniquely up to rescaling and changing coordinates.
-We now extend to SL operators (11) with $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq.
+We now extend to SL operators (11) with a smooth $q(x)\neq 0$, corresponding to an inhomogeneous (possibly imaginary) mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ by adding $v_{0}^{2}M(x)^{2}\varphi(x)^{2}$ inside the big parentheses in Eq.
 (8):
-$H_{\textrm{iTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\biggl{(}\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}+v_{0}^{2}M(x)^{2}\varphi(x)^{2}\biggr{)}\!:\right.\!,$ (49)
+$H_{\textrm{miTLL}}\equiv\int_{-L/2}^{L/2}\frac{\mathrm{d}x}{2\pi}v(x)\!\left.:\!\hskip-0.5pt\biggl(\frac{[\pi\Pi(x)]^{2}}{K(x)}+K(x)[\partial_{x}\varphi(x)]^{2}+v_{0}^{2}M(x)^{2}\varphi(x)^{2}\biggr)\hskip-0.5pt\!:\right.\!$.
+(61)
 Indeed, it is straightforward to extend the previous discussion to such models.
@@
 (13), thus $E_{0}>0$.
-Instead, we have a “massive” inhomogeneous TLL, for which one can show:
-A “massive” inhomogeneous TLL exhibits PWT if and only if $v(x)$, $K(x)$, $q(x)$ are even
-and all eigenvalues are of the form $E_{n}=\pi\tilde{m}_{n}/T$ for certain integers $\tilde{m}_{n}$.
-Here, these integers $\tilde{m}_{n}=m_{n}+c$ in terms of the integers $m_{n}$ for the massless case in the main text, where we allow for an overall shift by a finite even integer $c$.
-[The latter must be even so that $\tilde{m}_{n}$ is even (odd) if $n$ is even (odd), as required due to Property (ii) in the main text.] Moreover, given such a spectrum, the results in [82, 83] still imply that $\hat{q}(y)$ is uniquely determined by the spectrum, but this can no longer uniquely determine $K(x)$ since the first term in Eq.
-(18) is now non-zero in general.
+Instead, we have an inhomogeneous TLL with a mass term, for which one can show the following statement: “Massive” inhomogeneous TLLs exhibit PWT as defined in Eq.
+(1) if and only if $v(x)$, $K(x)$, and $q(x)$ are even and the eigenenergies are $E_{n}=\pi\tilde{m}_{n}/T$ for integers $\tilde{m}_{n}$ so that
+  1. (a)
+    $0\leq\tilde{m}_{0}<\tilde{m}_{1}<\ldots$ with each $\tilde{m}_{n}$ even or odd if $n$ is even or odd,
+  2. (b)
+    $\tilde{m}_{n}\sim n$ for large $n$, i.e., $\tilde{m}_{n}=n+o(n)$.
+The integers in the result above can be written as $\tilde{m}_{n}=m_{n}+k$ in terms of the integers $m_{n}$ for the massless case in the main text, allowing for an overall shift by a finite even integer $k$.
+[The latter must be even so that (a) above is satisfied, as required due to Property (ii) in the main text.] Now, given such a spectrum, the results in [105, 106] still imply that $\hat{q}(y)$ is uniquely determined by it, but this can no longer uniquely determine $K(x)$ since the first term on the right-hand side of Eq.
+(18) is nonzero in general.
 As an example, we revisit the case with $v(x)=v\sqrt{1-(2x/L)^{2}}$ and $K(x)=K\bigl[1-(2x/L)^{2}\bigr]^{\alpha}$ for $\alpha>-1/2$ and constants $v,K>0$ encountered in the main text, whose eigenfunctions were given by Gegenbauer polynomials $C^{(\alpha)}_{n}(2x/L)$, but now we include a fine-tuned potential $q(x)=-(2v\alpha/L)^{2}w(x)$, where we recall that $w(x)=K(x)/v(x)$.
@@
 (13) are $\lambda_{n}=(\pi v_{0}/L)(n+\alpha)^{2}$ with $v_{0}=2v/\pi$, meaning that $E_{n}=\pi(n+\alpha)v_{0}/L$.
-Consequently, any $c=\alpha=0,2,4,\ldots$ works, yielding a discrete family of models with fine-tuned inhomogeneous mass that exhibit PWT.
+Consequently, any $\alpha=k=0,2,4,\ldots$ works, yielding a discrete family of models with fine-tuned inhomogeneous mass that exhibit PWT.
 We stress that the mass in this example satisfies $M(x)\propto 1/L$ and therefore vanishes in the thermodynamic limit.
 Thus, one should not expect hallmarks of massive theories, such as exponential decay of two-point correlations with respect to the spatial separation.
+From the theory of regular SL operators and the Bohr-Sommerfeld/WKB quantization condition in the previous section, if one assumes regularity also after unfolding to $S^{1}_{2L}$, then “massive” inhomogeneous TLLs must lie in the orbit of $\hat{q}(y)=0$ in order to exhibit PWT.
+I.e., in that case, while $K(x)$ does not have to be constant, the choice of $q(x)$ must be such that $\hat{q}(y)$ given by Eq.
+(18) is still zero, corresponding to an exactly linear spectrum as stated in the main text.
+However, the above Gegenbauer example is irregular, meaning that it does not have to lie in the orbit of $\hat{q}(y)=0$ (which it does not for $\alpha=2,4,\ldots$) while exhibiting PWT.
 Liouville normal form and compactified free bosons
-It is curious to note the following “shortcut” to writing our massless bosons (8) in Liouville normal form: Recall that an inhomogeneous TLL is a theory of compactified free bosons with a position-dependent compactification radius $R(x)$ in curved spacetime given by the metric $(h_{\mu\nu})=\operatorname{diag}(v(x)^{2}/v_{0}^{2},-1)$ for coordinates $(x^{0},x^{1})\equiv(v_{0}t,x)$; see, e.g., [27, 53].
-Given a Luttinger parameter profile $K(x)$, the compactification radius is $R(x)=\sqrt{2\alpha^{\prime}}K(x)$ ($\alpha^{\prime}$ has dimension length squared and is commonly set equal to $2$).
+It is curious to note the following “shortcut” to writing our massless boson model (8) in Liouville normal form: Recall that inhomogeneous TLL theory is that of compactified free bosons in curved spacetime with a position-dependent compactification radius $R(x)$ and the metric $(h_{\mu\nu})=\operatorname{diag}(v(x)^{2}/v_{0}^{2},-1)$ in coordinates $(x^{0},x^{1})\equiv(v_{0}t,x)$; see, e.g., [37, 61].
+Given a Luttinger parameter profile $K(x)$, the compactification radius is $R(x)\equiv\sqrt{2\alpha^{\prime}}K(x)$ ($\alpha^{\prime}$ has dimensions of length squared and is commonly set equal to $2$).
 In terms of the bosonic field $\varphi(x)$ (with values taken modulo $2\pi$), the action corresponding to the Hamiltonian (8) is
-$S=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}R(x)^{2}(\partial_{\mu}\varphi)(\partial^{\mu}\varphi),$ (50)
+$S_{\textrm{iTLL}}\equiv\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}R(x)^{2}(\partial_{\mu}\varphi)(\partial^{\mu}\varphi),$ (62)
 where we recall that $h\equiv\det(h_{\mu\nu})$.
-This can be viewed as a non-exactly marginal deformation of a theory with constant compactification radius; cf.
+This can be viewed as a nonexactly marginal deformation of a theory with constant compactification radius; cf.
 [61].
 The corresponding action for $X(x)\equiv R(x)\varphi(x)$ becomes
-$S=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}\bigl{[}(\partial_{\mu}X)(\partial^{\mu}X)-\hat{q}X^{2}/v(x)^{2}\bigr{]}$ (51)
+$S_{\textrm{iTLL}}=\frac{1}{4\pi\alpha^{\prime}}\int\mathrm{d}^{2}x\,\sqrt{-h}\bigl[(\partial_{\mu}X)(\partial^{\mu}X)-\hat{q}X^{2}/v(x)^{2}\bigr]$ (63)
 with $\hat{q}=\hat{q}(y)$ in Eq.
@@
 (19) for $\mathfrak{u}(y)=X(x)/\sqrt{2\alpha^{\prime}}$.
-I.e., as a theory for $X(x)$, even when $q(x)=0$, the position-dependent $K(x)$ corresponds a (possibly imaginary) inhomogeneous mass $\sqrt{\hat{q}(y)}/v(x)$.
+I.e., as a theory in terms of the field $X(x)$, even when $q(x)=0$, the position-dependent $K(x)$ corresponds a (possibly imaginary) inhomogeneous mass $\sqrt{\hat{q}(y)}/v(x)$.
 This also extends to the case
-$S=\frac{1}{2\pi}\int\mathrm{d}^{2}x\sqrt{-h}\bigl{[}K(x)(\partial_{\mu}\varphi)(\partial^{\mu}\varphi)-v_{0}^{2}M(x)^{2}\varphi^{2}\bigr{]}$ (52)
-with a non-zero mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ for the bosonic field $\varphi(x)$, corresponding to the Hamiltonian (49), the only difference being that the first term in Eq.
+$S_{\textrm{miTLL}}\equiv\frac{1}{2\pi}\int\mathrm{d}^{2}x\sqrt{-h}\bigl[K(x)(\partial_{\mu}\varphi)(\partial^{\mu}\varphi)-v_{0}^{2}M(x)^{2}\varphi^{2}\bigr]$ (64)
+with a nonzero mass $M(x)=\sqrt{-q(x)/v_{0}^{2}v(x)}$ for the bosonic field $\varphi(x)$, corresponding to the Hamiltonian (61), the only difference being that the first term on the right-hand side of Eq.
 (18) is then no longer zero.
@@
 (8).
-Using the eigenfunctions obtained from SL theory, we can compute any ground-state two-point correlation functions for the fields $\varphi(x)$ and $\vartheta(x)$, where the latter is defined by
+Using the eigenfunctions obtained from SL theory, we can compute any correlation function for the fields $\varphi(x)$ and $\vartheta(x)$, where the latter is defined by
 $\partial_{x}\vartheta(x)\equiv\pi\Pi(x).$ (65)
-Indeed, using the expansions (45), that $a_{n}$ annihilates the ground state $|\Omega\rangle$ of $H_{\textrm{iTLL}}$, as well as the commutation relations $\bigl{[}a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr{]}=\delta_{n,n^{\prime}}$, it is straightforward to show that
+Case in point, using the expansions in Eq.
+(45) with zero modes omitted, that $a_{n}$ annihilates the ground state $|\Omega\rangle$ of $H_{\textrm{iTLL}}$, and the commutation relations $\bigl[a^{\vphantom{\dagger}}_{n^{\vphantom{\prime}}},a^{\dagger}_{n^{\prime}}\bigr]=\delta_{n,n^{\prime}}$, it is straightforward to show that the ground-state two-point correlations are
 $\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ $=\sum_{n=1}^{\infty}\frac{\pi}{2E_{n}}u_{n}(x)u_{n}(x^{\prime})\mathrm{e}^{-\mathrm{i}E_{n}(t-t^{\prime})},$ (66)
@@
 $\partial_{x}U_{n}(x)\equiv w(x)u_{n}(x).$ (67)
-Similar correlation functions can be computed for other objects, including vertex operators (i.e., normal-ordered exponentials of the bosonic fields) [54, 55, 56, 57, 58, 59].
-It is also straightforward to compute Green’s functions; see, e.g., [27], the only difference with the present case being that one must account for the non-trivial weight function $w(x)$ that was simply a constant in [27].
+The first of these correlations is the same as the one in Eq.
+(17) in the main text.
+Similar correlations can be computed for other objects, including vertex operators, i.e., normal-ordered exponentials of the bosonic fields [62, 63, 64, 65, 66, 67].
+It is also straightforward to compute Green’s functions; see, e.g., [37], the only difference with the present case being that one must account for the nontrivial weight $w(x)$ that was simply a constant in [37].
 As an important example, consider a theory with $K(x)=K>0$ constant but $v(x)>0$ general [and $q(x)=0$].
-We recall that the SL operator in this case is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using Eq.
+We recall that the SL operator in this case is $\mathcal{A}=-v(x)\partial_{x}v(x)\partial_{x}=-v_{0}^{2}\partial_{y}^{2}$ using $y=y(x)$ given by Eq.
 (7) with eigenenergies $E_{n}=\pi nv_{0}/L$ and eigenfunctions $u_{n>0}(x)=\sqrt{2v_{0}/L}\cos\bigl(n\pi[L+2y]/2L\bigr)$ and $u_{0}(x)=\sqrt{v_{0}/L}$.
-Then, one can show that
-$\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle\\ \begin{aligned} =\frac{1}{4}&\left(\log\frac{\mathrm{i}\exp\Bigl{(}-\mathrm{i}\frac{\pi}{2L}[f(x)-f(x^{\prime})-v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)-f(x^{\prime})-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr{)}}+\log\frac{\mathrm{i}\exp\Bigl{(}-\mathrm{i}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L-v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr{)}}\right.\\ &\quad+\left.\log\frac{-\mathrm{i}\exp\Bigl{(}\mathrm{i}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L+v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)+f(x^{\prime})+L+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr{)}}+\log\frac{-\mathrm{i}\exp\Bigl{(}\mathrm{i}\frac{\pi}{2L}[f(x)-f(x^{\prime})+v_{0}(t-t^{\prime})]\Bigr{)}}{2\sin\Bigl{(}\frac{\pi}{2L}[f(x)-f(x^{\prime})+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr{)}}\right),\end{aligned}$ (56)
-where we included the necessary regularizations in order to interpret coincident spacetime points and used the identity $\sum_{n=1}^{\infty}\frac{1}{n}\mathrm{e}^{-\xi n}=-\log\left(1-\mathrm{e}^{-\xi}\right)$ for any $\xi$ such that $\operatorname{Re}\xi>0$.
-Eq.
-(56) is the analytical expression for what was plotted in Fig.
+One can then show that the expression (17) for the ground-state $\varphi\varphi$ correlation function (with zero modes omitted) yields
+$\hskip-6.25958pt\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle\\ \begin{aligned} =\frac{1}{4}&\left(\log\frac{\mathrm{i}\exp\Bigl(-\mathrm{i}\frac{\pi}{2L}[y(x)-y(x^{\prime})-v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)-y(x^{\prime})-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr)}+\log\frac{\mathrm{i}\exp\Bigl(-\mathrm{i}\frac{\pi}{2L}[y(x)+y(x^{\prime})+L-v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)+y(x^{\prime})+L-v_{0}(t-t^{\prime})+\mathrm{i}0^{+}]\Bigr)}\right.\\ &\quad+\left.\log\frac{-\mathrm{i}\exp\Bigl(\mathrm{i}\frac{\pi}{2L}[y(x)+y(x^{\prime})+L+v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)+y(x^{\prime})+L+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr)}+\log\frac{-\mathrm{i}\exp\Bigl(\mathrm{i}\frac{\pi}{2L}[y(x)-y(x^{\prime})+v_{0}(t-t^{\prime})]\Bigr)}{2\sin\Bigl(\frac{\pi}{2L}[y(x)-y(x^{\prime})+v_{0}(t-t^{\prime})-\mathrm{i}0^{+}]\Bigr)}\right),\end{aligned}$ (68)
+where we included necessary regularizations in order to interpret coincident spacetime points and used the identity $\sum_{n=1}^{\infty}\frac{1}{n}\mathrm{e}^{-\xi n}=-\log\left(1-\mathrm{e}^{-\xi}\right)$ for any $\xi$ such that $\operatorname{Re}\xi>0$.
+The real part of Eq.
+(68) gives the analytical expression for what was plotted in Fig.
 1 (assuming $K=1$).
-Moreover, for delta-like waves, the terms appearing in $\log F(t)$ given by Eq.
+Moreover, for deltalike waves, the terms appearing in $\log F(t)$ given by Eq.
 (41) for the free-fermion case (37) closely resemble those in Eq.
 (68).
-Indeed, the former agrees with $4\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ to leading order, and one can furthermore show that the agreement between the two ways of computing (CFT methods versus bosonization) is exact (as it should be) by evaluating and summing all combinations of two-point correlation functions of the vertex operators describing the right/left-moving free fermionic fields.
+Indeed, the former agrees with $4\langle\Omega|\varphi(x,t)\varphi(x^{\prime},t^{\prime})|\Omega\rangle$ to leading order, and one can furthermore show that the agreement between the two ways of computing (CFT methods versus bosonization) is exact (as it should be) by evaluating and summing all combinations of two-point correlation functions of the vertex operators describing the right- and left-moving free fermionic fields.